### v2

# Privacy Guarantees in Posterior Sampling under Contamination

Abstract
In recent years differential privacy has been adopted by tech-companies and governmental agencies as the standard for measuring privacy in algorithms. In this article, we study differential privacy in Bayesian posterior sampling settings. We begin by considering differential privacy in the most common privatization setting in which Laplace or Gaussian noise is simply injected into the output. In an effort to achieve better differential privacy, we consider adopting Huber’s contamination model for use within privacy settings, and replace at random data points with samples from a heavy-tailed distribution (instead of injecting noise into the output). We derive bounds for the differential privacy level $(\epsilon,\delta)$ of our approach, without the need to impose the restriction of having a bounded observation and parameter space which is commonly used by existing approaches and literature. We further consider for our approach the effect of sample size on the privacy level and the convergence rate of $(\epsilon,\delta)$ to zero. Asymptotically, our contamination approach is fully private at no cost of information loss. We also provide some examples depicting inference models that our setup is applicable to with a theoretical estimation of the convergence rate, together with some simulations.
62F15,
62J12,
Huber’s Contamination Model,
Posterior Sampling,
Bayesian Inference,
Differential Privacy,
keywords:
[class=MSC]
keywords:
[…]
, , , , and
1 Introduction
With the rapid growth of smart electronic devices in recent decades, the quantity of data generated and collected every day has increased by orders of magnitude. A byproduct of this expansion has led to increased awareness of personal privacy and concerns about privacy leakage as a consequence of utilizing the collected data in studies. In recent years, differential privacy has become of significant interest to computer scientists, statisticians and others for studying the properties and trade-offs in designing algorithms that protect the privacy of individuals, and also being increasingly adopted by tech-companies, for instance, Google (Erlingsson, Pihur and Korolova, 2014; Aktay et al., 2020), Microsoft (Ding, Kulkarni and Yekhanin, 2017; Pereira et al., 2021), Apple (Apple, 2017), or government agencies (Machanavajjhala et al., 2008; Abowd, 2018).
Differential privacy, as defined in Dwork et al. (2006, 2014), asserts the difference in response of a randomized algorithm $\mathcal{A}:\mathcal{X}^{n}\rightarrow\Omega$ when being supplied two neighbouring datasets of size $n$, i.e., one dataset is a modification of the other dataset by changing only one data point. In particular, the response pattern should satisfy the following inequality for any measurable set $S\subset\Omega$ and any neighbouring datasets $\bm{X}$ and $\bm{Z}\in\mathcal{X}^{n}$,
$\mathbb{P}(\mathcal{A}(\bm{X})\in S)\leq\exp(\epsilon)\mathbb{P}(\mathcal{A}(\bm{Z})\in S)+\delta.$ (1)
For simplicity, in this paper, we consider $\mathcal{X}$ to be $\mathbb{R}^{d}$ or a convex subset of $\mathbb{R}^{d}$. The algorithm is considered ”more private” when $\epsilon$ and $\delta$ are smaller, in the sense that an arbitrarily large change in a single data point does not incur a significant change in the response pattern.
1.1 Related Literature
In most cases, differential privacy is achieved by purposefully injecting noise into the computation process which perturbs the final outcome. The injection of noise can happen in two places, either at the individual level before data collection or after data collection before releasing the result. The most studied mechanism for privatizing Bayesian inference is the latter, where the outcome is perturbed for privacy. Mainly, there are two types of problems being targeted: direct estimation problems (Foulds et al., 2016; Wang, Fienberg and Smola, 2015; Zhang, Rubinstein and Dimitrakakis, 2016; Bernstein and Sheldon, 2019; Kulkarni et al., 2021) and sampling problems (Dimitrakakis et al., 2013; Zhang, Rubinstein and Dimitrakakis, 2016; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019).
The typical approach for privatizing estimation problems is to directly inject Laplacian noise into the steps or the output (Wang, Fienberg and Smola, 2015; Foulds et al., 2016; Zhang, Rubinstein and Dimitrakakis, 2016). Bernstein and Sheldon (2019); Kulkarni et al. (2021) considered the injection of noise after moment approximations of the sufficient statistics to achieve privacy. Gaussian/Laplacian mechanisms suffer from the limitation that the bandwidth of the noise scales with the sensitivity, defined as the maximal difference in response with respect to a pair of neighbouring datasets of the target function. In most applications, the sensitivity is not bounded unless the observation space is bounded.
It is worth noting that the definition of ”Bayesian Differential Privacy” in Triastcyn and Faltings (2019, 2020) describes a different problem compared to the previously mentioned papers studying differential privacy in Bayesian inference following Dwork et al. (2006)’s definition. Triastcyn and Faltings (2019, 2020)’s work treats the dissimilar data point $z$ as a random variable, which is marginalized, as opposed to in (1) where both $\bm{X}$ and $\bm{Z}$ are considered as arbitrary but fixed.
In this paper, we consider the differential privacy property of Bayesian posterior sampling problem $\theta\sim\pi_{n}(\cdot|\bm{X})$ where $\pi_{n}$ is the posterior distribution given the dataset $\bm{X}\in\mathcal{X}^{n}$ of size $n$, modelled by likelihood function $f(x;\theta)$, $\theta\in\Omega$ and prior $\pi_{0}(\theta)$. Thereafter, we will call $\mathcal{X}$ the observation (or data) space where the input data resides and $\Omega$ the parameter space where the model parameter $\theta$ takes values in.
As noted in Dimitrakakis et al. (2013), Bayesian posterior sampling is itself a differentially private mechanism without additional noise under the condition that the log density is Lipschitz continuous in data with respect to some pseudo metric $\rho$ for any fixed parameter $\theta$. The level of privacy can be quantified by the pseudo distance $\rho$ between the dissimilar data points $\rho(\bm{x},\bm{z})$. However, allowing the level of privacy to depend on $\rho(\bm{x},\bm{z})$ implies the algorithm can be arbitrarily non-private when the observation space is unbounded. This limitation is shared by the line of studies following Dimitrakakis et al. (2013) on posterior sampling due to having a similar setup (Wang, Fienberg and Smola, 2015; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019).
An additional difficulty in bounding the privacy level of posterior sampling comes from the dependency of sensitivity on the parameter space. The sensitivity of the posterior sampling problem is determined by the sensitivity of the log-posterior density function which depends directly on the sample $\theta$. Thus, the sensitivity could also be unbounded if the parameter space is unbounded (Zhang, Rubinstein and Dimitrakakis, 2016; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019; Zhang and Zhang, 2023).
In contrast to adding noise at the outcome level, we consider Huber’s contamination model (Huber, 2004) as a differential privacy mechanism. A body of literature exists in quantifying the trade-off between privacy and efficiency statistical inference under Huber’s contamination model (Rohde and Steinberger, 2020; Cai, Wang and Zhang, 2021; Kroll, 2021; Li, Berrett and Yu, 2022). However, these analyses are local differential privacy models, where the privacy is assessed with respect to datasets of size $1$, whereas we study differential privacy of the whole inference/sampling algorithm under Huber’s contamination model, otherwise known as global differential privacy model.
1.2 Our contributions
The novel contributions of this work are as follows:
  • We derive the differential privacy properties of a general posterior sampling problem under Huber’s contamination model (Huber, 2004) without assuming bounded parameter space or observation space, which are conditions that often appear explicitly or implicitly in the existing literature to ensure finite sensitivity;
  • We characterise the rate of convergence for the differential privacy cost $(\epsilon,\delta)$ with the number of data points $n$ which, to the best of our knowledge, is yet to be answered;
  • We also show through simulations the method is more robust since the injected noise is independent of the size of the target space (Section 2.3), and introducing some contamination could vastly improve the differential privacy when the original model is differentially private by design (Section 4.2).
Let $k_{p}$ denote the contaminated likelihood with contamination rate $p$, written as a mixture of the original likelihood function $f$ and a contamination density $g$,
$k_{p}(\bm{x};\theta):=(1-p)f(\bm{x};\theta)+pg(x).$
Under a set of mild assumptions on the prior and likelihood function defined later in Section 3.2, we informally state our main result:
Theorem 1.
Let $\bm{X}$,$\bm{Z}$ be neighbouring datasets of size $n$ generated by $\mathbb{P}_{\theta^{*}}$, the true model under contamination, with density function $k_{p_{n}}(\cdot;\theta^{*})$. Under mild assumptions on the set of density functions $k_{p}$ and prior $\pi_{0}$ (Assumptions 1-5 defined in Section 3.2), then there exists a sequence of contamination probabilities $\{p_{n}\}\rightarrow 0$, such that $\forall S$ measurable, $\epsilon,\delta>0$, we have
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp(\epsilon)\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta,$
$\mathbb{P}_{\theta^{*}}$-almost surelly, where $\mathbb{P}_{\pi_{n}}(S|\bm{X})$ denotes the probability mass of $S$ under posterior $\pi_{n}$ given dataset $\bm{X}$. In addition, the entries of Fisher information also converge to the true model
$I_{p_{n}}(\theta^{*})\rightarrow I(\theta^{*}).$
where $I_{p_{n}}$ is the Fisher information matrix for the contaminated likelihood when the contamination rate is $p_{n}$ and $I(\cdot)=I_{0}(\cdot)$.
Theorem 1 essentially states that, under mild conditions, as the size of the dataset $n\rightarrow\infty$, sampling from the contaminated posterior achieves perfect differential privacy. Additionally, no statistical efficiency is lost since the contamination rate $p_{n}\rightarrow 0$.
1.3 Organization of the Paper
This paper is organized as follows. In Section 2 we introduce the setup of differential privacy and data contamination mechanism, and we demonstrate in Section 2.3 how our approach could improve in statistical efficiency compared with the standard frequentists’ approach to differential privacy in parameter estimation.
Following that, we present our main results on the privacy level of $p$-contaminated posterior sampling in Section 3. Any intermediate results for the main theorems are presented in Appendix A with the longer proofs in the supplementary material (Hu et al., 2024). We aim to lift both restrictions on having bounded parameter space and bounded observation space and derive a bound for privacy level that shrinks with $n$ in Section 3.2, with some discussions and remarks on the assumptions.
Since the link with the $(\epsilon,\delta)$ level and the choice of contamination $p$, $g(\cdot)$ is implicit. To help practitioners tune the model, we present, in Section 4, a simulation-based approach to empirically estimate $\epsilon,\delta$ one gets from a certain model setup, and empirically show that the trace of $(\epsilon,\delta)$ against $n$ follows what is described in Theorem 5 through simulation. On the theory side, we show in Section 5 and in the supplementary material (Appendix C) that our assumptions can be easily satisfied for many regression models when $n$ is large enough. Finally, we conclude the paper with a short discussion in Section 6.
2 Differential Privacy in Bayesian Sampling and Inference
In this section, we first introduce the setup of differential privacy in Bayesian posterior sampling and our contamination mechanism, followed by a comparison with the common frequentist’s approaches to privatising parameter estimation.
2.1 Differential Privacy in Bayesian Inference
In a typical Bayesian inference setting, we have access to a dataset of size $n$, $\mathcal{D}\in(\mathcal{X})^{n}$ where $\mathcal{X}$ is the space of all possible observations, modelled by a density function $f(\bm{x};\theta):\mathcal{X}\times\Omega\rightarrow\mathbb{R}_{\geq 0}$ with parameter $\theta$. The parameter $\theta$ is a random variable with prior density $\pi_{0}(\theta)$ defined on the measurable space $(\Omega,\mathcal{B}(\Omega))$. Then the posterior distribution for $\theta$ given the dataset $\mathcal{D}$ can be expressed as
$\pi_{n}(\theta|\mathcal{D})=\frac{\prod_{i=1}^{n}f(\bm{x}_{i};\theta)\pi_{0}(\theta)}{\int_{\Omega}\prod_{i=1}^{n}f(\bm{x}_{i};\theta)\pi_{0}(\theta)\mathtt{d}\theta}.$ (2)
Throughout this paper, we assume the following:
  • The likelihood functions $f(\bm{x};\theta)$ and $k_{p}(\bm{x};\theta)$ are jointly continuous in $x$ and $\theta$, and the contamination likelihood $g(x)$ has full support over the observation space;
  • The observation space $\mathcal{X}$ and the parameter space $\Omega$ are both convex subsets of the Euclidean space, but not necessarily bounded or of the same dimension.
For now, we also assume that the model is weakly consistent at $\theta^{*}$ (e.g., Definition 6.1 Ghosal and Van der Vaart (2017)). In the later sections, this weak consistency assumption will be implied by other assumptions.
Definition 1 ($(\epsilon,\delta)$-Differentially Private Posterior Sampling).
The direct sampling from the posterior distribution in (2) is considered to be $(\epsilon,\delta)$-differentially private if for any two datasets $\bm{X},\bm{Z}\in\mathcal{X}^{n}$ which differ by at most one entry, for all $S\in\mathcal{B}(\Omega)$,
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp({\epsilon})\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta.$ (3)
Definition 2 (Neighbouring Datasets).
Two datasets $\bm{X}$ and $\bm{Z}$ are neighbours to each other if and only if $\exists\mathcal{D}$ and $\bm{x},\bm{z}\notin\mathcal{D}$ such that $\bm{X}=\mathcal{D}\cup\{\bm{x}\}$, $\bm{Z}=\mathcal{D}\cup\{\bm{z}\}$.
Remark 1.
Definition 2 is perhaps the most adopted definition for neighbouring datasets in the differential privacy literature where the two datasets are explicitly required to have the same cardinality.
In some places, it is alternatively defined as one dataset is a proper subset of the other with strictly one less element, for instance, in Dwork (2008). To avoid ambiguity, we call two datasets $\bm{X}$ and $\bm{Z}$ strictly adjacent if $\bm{Z}=\bm{X}\cup\{\bm{z}\}$ and $\bm{z}\notin\bm{X}$ (or the same statement with $\bm{Z}$ and $\bm{X}$ swapped holds). Our result transports easily to the strictly adjacent case which is discussed in Appendix D in the supplementary material.
Remark 2.
In general, Bayesian posterior sampling is not differentially private. For instance, modelling the likelihood using a Gaussian distribution with known variance $1$ and a Gaussian prior $\mathcal{N}(\mu_{0},1)$ on the mean $\mu$. Let $\bm{X},\bm{Z}$ be neighbouring datasets with $\mathcal{D}=\bm{X}\cap\bm{Z}$, $\bm{X}=\mathcal{D}\cup\{x\}$, $\bm{Z}=\mathcal{D}\cup\{z\}$, where $\mathcal{D}:=\{x_{1},\dots,x_{n}\}$. Then
$\log\left(\frac{\pi(\mu|\bm{X})}{\pi(\mu|\bm{Z})}\right)=-\frac{1}{2(n+2)}(x^{2}-z^{2})+(x-z)\left(\mu-\frac{1}{n+2}\sum_{i=1}^{n}(x_{i}+\mu_{0})\right),$
which is unbounded in $\mu$, $x$ and $z$.
2.2 p-Contaminated Data
To improve privacy, usually, some noise is injected into the process. For instance, noise can be injected directly into the posterior distribution to contaminate the output sample. Here, we consider the injection of noise before the inference stage where for any data point $y$ in the dataset $\mathcal{D}$, we replace it with $\mathcal{C}(y)$ by
$\mathbb{P}(\mathcal{C}(\bm{y})=\bm{y})$ $\,=1-p,$
$\mathbb{P}(\mathcal{C}(\bm{y})=\bm{x})$ $\,=p,$
where $10<p<$ 1, $\bm{x}\sim g(\cdot)$ and $g(\cdot)$ is the density function which has a heavier tail than $f(\bm{x};\theta),\forall\theta$. Thus, the likelihood function of the contaminated dataset is given by $k_{p}(\bm{x};\theta)=(1-p)f(\bm{x};\theta)+pg(\bm{x})$.
Remark 3.
We may consider two distinct DP setups with the same contamination strategy. When we discuss the differential privacy of a randomized algorithm $\mathcal{A}$ (see (1), it is important to identify what random process is involved in $\mathcal{A}$, in this case, whether contamination is part of $\mathcal{A}$, i.e., contamination at runtime, or not part of $\mathcal{A}$, i.e., contamination before $\mathcal{A}$. In our setup, the latter is preferred, due to its simplicity. Though, one may also consider the former approach where the uncontaminated neighbouring datasets $\bm{X}$ and $\bm{Z}$ are contaminated at runtime to produce $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$. Now, $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$ are random variables and their realizations don’t need to be neighbouring to each other. It is not hard to show that if the pre-contaminated sampling algorithm is $(\epsilon,\delta)$-DP, then dynamic contamination will be at least $(\epsilon,\delta)$-DP. However, since at runtime, $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$ are not necessarily neighbouring data sets, the resulting problem is more complicated to analyze. Either way, one should treat the contamination density $g(\cdot)$ and the contamination rate $p$ as public knowledge to avoid having model identifiability issues (Mu and Xiong, 2023).
Hereafter, we will assume the dataset to be precontaminated and denote $\bm{X}$ as the result of the contamination. Thus the hypothetical neighbouring dataset $\bm{Z}$ is a neighbour to the already contaminated dataset $\bm{X}$ and thus $\bm{Z}$, obtained from arbitrarily changing one entry of $\bm{X}$, contains the same contamination as $\bm{X}$ (possibly except for one entry).
2.3 Comparing frequentist and Bayesian approaches in DP
Before we introduce the main result, we present a motivating example for seeking a differential privacy bound that does not require assumptions of bounded parameter or observation space. The most common frequentist approach to retaining privacy is to add noise directly to the output (Dwork et al., 2006, 2014). Clearly such an approach will (like our contamination methodology) lose statistical efficiency. Within this simple example, we can directly compare the loss of efficiency that results from insisting on particular levels of DP.
The level of noise added by frequentist approaches usually grows with the size of the bounded space. Thus, this example shows how in the finite sample regime, the Bayesian approach can yield better statistical efficiency than the frequentist approaches especially with small datasets, even though asymptotically no Bayesian mechanism can beat the optimal frequentist estimator, as recognised in Wang, Fienberg and Smola (2015), due to Bernstein-von Mises theorem (Van der Vaart, 2000).
Consider the following simple linear regression model with no covariates:
$x_{i}=\theta+e_{i}$
where the true value $\theta^{*}$ is assumed to be $30$, and $e_{i}\sim\mathcal{N}_{[l,u]}(0,\sigma^{2})$ are i.i.d. truncated normal distributions supported on interval $[l,u]$, with standard deviation $\sigma=8$.
Let $\bm{X}:=\{x_{1},\dots,x_{n}\}$ be the set of observations, for given contamination rate $p_{n}:=\exp(-n^{1/12})$, the dataset is contaminated by replacing with probability $p_{n}$ each observation using $\tilde{x}_{i}\sim T_{\nu}(\theta^{*},\sigma)$ supported on $[l,u]$. The density of the scaled and truncated t-distribution is given by
$g(\tilde{x}_{i})\propto\left(1+\frac{(\tilde{x}_{i}-\theta^{*})^{2}}{\nu\sigma^{2}}\right)^{-\tfrac{\nu+1}{2}}\mathbb{I}_{l\leq\tilde{x}_{i}\leq u},\quad\nu=5.$
Then we empirically derive $\hat{\epsilon}$ and $\hat{\delta}$ with respect to the dataset by using the approach we will describe later in Section 4. Finally, we produce the following three different estimations of $\theta$ with respect to the same privacy level $(\hat{\epsilon},\hat{\delta})$:
  1. 1.
    Bayesian: draw $\hat{\theta}_{B}\sim\pi_{n}(\cdot|\bm{X})$;
  2. 2.
    Frequentist (Laplace noise; Dwork et al. (2006)):
    $\hat{\theta}_{L}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}+\text{Laplace}(\tfrac{u-l}{n\hat{\epsilon}});$ (4)
  3. 3.
    Frequentist (Guassian noise; Dwork et al. (2014)):
    $\hat{\theta}_{N}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}+\text{Normal}\left(0,2\log\tfrac{1.25}{\hat{\delta}}\left(\tfrac{u-l}{n\hat{\epsilon}}\right)^{2}\right).$ (5)
Then by definition, $\hat{\theta}_{B}$ and $\hat{\theta}_{N}$ are $(\hat{\epsilon},\hat{\delta})$-DP and $\hat{\theta}_{L}$ is $(\hat{\epsilon},0)$-DP.
Since the implementation of Laplace or Gaussian noise requires a boundedness assumption observations $x_{i}\in[l,u]$, $\forall i$. To demonstrate the effect of changing the interval size, the simulation is conducted on two different support intervals $[l,u]=[0,300]$ and $[l,u]=[0,600]$. For both support sizes, the simulation is repeated 50 times, with respect to the same dataset $\bm{X}$, and the empirical variance of the estimators is computed. The results are summarized in Tables 0(a) and 0(b) respectively.
Since the frequentist approaches are unbiased, but the Bayesian approach is biased, we use the Mean Squared Error (MSE) as a measure of statistical efficiency. Since both frequentist estimators are unbiased, the MSE is dominated by the variance of the injected noise. Noticeably, both (frequentist) estimators roughly quadrupled in variance as the interval size doubled, which matches (4) and (5). On the other hand, we notice that the Bayesian estimator is mostly unaffected, both in empirical estimates of $\hat{\epsilon}$, $\hat{\delta}$ and in the MSE, when the interval size doubled.
In addition, the contamination probabilities are roughly $0.23$ and $0.17$ when $n=100$ and $1000$ respectively. Nonetheless, in these cases, the Bayesian approach still offers a better trade-off between statistical efficiency and privacy since the estimator has far better MSE compared with the two frequentist approaches.
This shows that while the frequentist’s estimator can achieve the optimal variance when the Cramer-Rao lower bound is attained in the asymptotic regime, a conservative bound assumption can lead to unnecessarily large statistical noise, especially when the dataset is moderately small. In contrast, the Bayesian approach gives a more robust estimator, in the sense that the level of differential privacy and estimator variance it achieves is independent of the size of the underlying space.
Table 1: Mean Square Error Comparison for Frequentist and Bayesian Estimators
$n$ $\hat{\epsilon}$ $\hat{\delta}$ MSEL MSEG MSEB
100 0.34 0.61 164.67 133.52 4.48
1000 0.23 0.22 3.34 6.54 0.33
5000 0.19 0.027 0.24 0.95 0.07
10000 0.16 0.003 0.07 0.47 0.03
(a) $l=0$, $u=300$
$n$ $\hat{\epsilon}$ $\hat{\delta}$ MSEL MSEG MSEB
100 0.34 0.61 658.08 537.81 4.30
1000 0.23 0.22 12.45 26.15 0.35
5000 0.19 0.03 0.87 3.73 0.07
10000 0.17 0.003 0.27 1.71 0.03
(b) $l=0$, $u=600$
2.4 Posterior Decomposition
Here we quickly outline our strategy to bound $\epsilon$ and $\delta$. Let
$d(\bm{x},\theta):=\frac{k_{p}(\bm{x},\theta)}{k_{p}(\bm{x},\theta^{*})}$
where $\theta^{*}$ is the true parameter. If $X$ is a dataset with $n$ data points $x_{i},i=1,\dots,n$, then for $A\subset\Omega$, denote
$m_{n}(A,\bm{X})=\int_{A}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\theta)\mathtt{d}\theta.$
Now, we can write the posterior distribution function $\pi_{n}(\cdot|\bm{X})$ in terms of $m_{n}(\cdot,\bm{X})$,
$\mathbb{P}_{\pi_{n}}(A|\bm{X})=\frac{\int_{A}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}=\frac{m_{n}(A,\bm{X})}{m_{n}(\Omega,\bm{X})}.$ (6)
Let $A_{n}:=\{\theta\in\Omega:h(\theta,\theta^{*})\leq\phi_{n}\}$, defined by a positive sequence $\{\phi_{n},n\in{\mathbb{N}}\}\rightarrow 0$ and $h$ is the Hellinger distance defined (later) in Definition 3. $A_{n}$ may be defined using other distance metrics such that $\mathbb{P}_{\pi_{n}}(A_{n}^{c})\rightarrow 0$. We will prove our main theorems for Hellinger distance but use $L_{\infty}$ distance in the simulation. Then
$\mathbb{P}_{\pi_{n}}(S|\bm{X})=\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})+\mathbb{P}_{\pi_{n}}(S\cap A_{n}^{c}|\bm{X})$
where we aim to bound $\mathbb{P}_{\pi_{n}}(S\cap A_{n}^{c}|\bm{X})$ by $\delta$, so that $\epsilon$ is only required to be upper bounded within a compact set $A_{n}$.
Proposition 2.
Let $\bm{X}$ and $\bm{Z}$ be neighbouring datasets, i.e., $\bm{X}=\mathcal{D}\cup\{\bm{x}\}$, $\bm{Z}=\mathcal{D}\cup\{\bm{z}\}$, $\bm{x},\bm{z}\notin\mathcal{D}$. Let $A_{n}$ be defined as above, then for any $S\subset\Omega$,
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\eta_{n}\left[\eta_{n}+\frac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$ (7)
where
$\eta_{n}=\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}.$ (8)
Now if we can find the upper bounds for $m_{n}(A_{n}^{c},\bm{X})$, $m_{n}(A_{n}^{c},\bm{Z})$ and $\eta_{n}$, and the lower bound for $m_{n}(\Omega,\bm{X})$ to upper bound the above expression, then we can treat (the upper bounds of) $\eta_{n}(\eta_{n}+\tfrac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})})$ as $e^{\epsilon}$ and $\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$ as $\delta$.
Proof for Proposition 2.
Note that using the notation of (6),
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}^{c}|\bm{X})\leq\mathbb{P}_{\pi_{n}}(A_{n}^{c}|\bm{X})=\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$
and
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{\int_{A_{n}\cap S}\prod_{i=1}^{n}d(\bm{z}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}$
$\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{\int_{\Omega}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}{\int_{\Omega}\prod_{i=1}^{n}d(\bm{z}_{i},\theta)\pi_{0}(\mathtt{d}\theta)}\mathbb{P}_{\pi_{n}}(A_{n}\cap S|\bm{Z})$
$\leq$ $\,\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{m_{n}(A_{n},\bm{Z})+m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\mathbb{P}_{\pi_{n}}(A_{n}\cap S|\bm{Z})$ (9)
3 Generic Results on DP Guarantees under Bayesian Sampling
This section will be devoted to presenting our main results and discussing the required assumptions.
3.1 Preliminaries
When the parameter space is non-compact, we need some control over the space complexity of density functions to compute the tail bound for the posterior distribution, which will act as an upper bound of $\delta$. In empirical processes, covering numbers and bracketing numbers are often used to quantify the size and complexity of a metric space.
Definition 3 (Hellinger Distance).
Let $p,q$ be two density functions defined on a measurable space $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, then the (squared) Hellinger distance between $p$ and $q$ is defined as
$h^{2}(p,q):=\frac{1}{2}\left\|p^{1/2}-q^{1/2}\right\|_{2}^{2}=\frac{1}{2}\int_{\mathbb{R}^{d}}\left(\sqrt{p(\bm{x})}-\sqrt{q(\bm{x})}\right)^{2}\mathtt{d}\bm{x}.$
Throughout this paper, the Hellinger distance will be used as the preferred metric for computing bracketing numbers, unless specifically stated. We hereafter omit the inclusion of metric choice in the notations.
When the context is clear, we will use $h(\theta_{1},\theta_{2})$ to denote the distance between two contaminated density functions $h(k_{p}(\cdot;\theta_{1}),k_{p}(\cdot;\theta_{2}))$, and replace the function space using the parameter space $\Omega$.
Let $\Xi$ denote the set of non-negative integrable functions on $\mathbb{R}^{d}$ and $\mathcal{P}\subset\Xi$ be the set of density functions of our interest.
Definition 4 ($r$-Hellinger Bracketing).
A set of function pairs $\{[L_{j},U_{j}],j=1,\dots,N\}\subset\Xi\times\Xi$ is a $r$-Hellinger bracketing of the space $\mathcal{P}$, if
$h(L_{j},U_{j})\leq\epsilon\text{ for all }j\in\{1,\dots,N\}$
and
$\forall p\in\mathcal{P},\,\,\exists j\in\{1,\dots,N\}\text{ such that }L_{j}(\bm{x})\leq p(\bm{x})\leq U_{j}(\bm{x}),\forall\bm{x}\in\mathbb{R}^{d}.$
An $\epsilon$-bracket is a pair $[L_{j},U_{j}]$ with $h(L_{j},U_{j})\leq r$.
Remark 4.
The Hellinger distance is a metric on density functions, but it also generalizes easily to a metric on the space of non-negative integrable functions for the purpose of the above definition.
Definition 5 (Hellinger Bracketing Number).
The $\epsilon$-Hellinger bracketing number is size of the smallest $\epsilon$-bracketing, defined as
$N_{[]}(\epsilon,\mathcal{P}):=\min\{|B|:B\text{ is an $\epsilon$-Hellinger bracketing of $\mathcal{P}$}\}.$
The $\bm{\epsilon}$-Hellinger bracketing entropy is the log of the bracketing number
$H_{[]}(\epsilon,\mathcal{P}):=\log N_{[]}(\epsilon,\mathcal{P}).$
3.2 Main Results
First, we define some necessary assumptions.
For $A\subset\Omega$, denote the set of density functions $f(\bm{x};\theta)$ indexed by $\theta\in A$ by $\mathcal{F}(A):=\{f(\cdot;\theta):\theta\in A\}$. Similarly, $\mathcal{K}_{p}(A):=\{k_{p}(\cdot;\theta);\theta\in A\}$.
Assumption 1.
There exists a positive sequence $\phi_{n}\rightarrow 0$ with $n\phi_{n}^{2}\rightarrow\infty$, constants $c>0$ with a sequence of expanding subsets $\Omega_{1}\subset\Omega_{2}\subset\dots\subseteq\Omega$, $\Omega_{n}\rightarrow\Omega$, such that
$\int_{\phi_{n}^{2}/2^{10}}^{\phi_{n}}H^{1/2}_{[]}(u,\mathcal{F}(\Omega_{n}))\mathtt{d}u\leq c\sqrt{n}\phi_{n}^{2}.$ (10)
Assumption 1 measures the space complexity of the density functions $f(\cdot;\theta)$. Note that $H_{[]}(u,\mathcal{F}(\Omega_{n}))\leq c^{2}n\phi_{n}^{2},\forall u\in(\phi_{n}^{2}/2^{10},\phi_{n})$ implies (10), then this condition can be verified easily with the following lemma.
Lemma 3.
Let $\Omega_{1}\subset\Omega_{2}\subset\dots\subset\Omega$ be a sequence of subsets such that $\Omega_{n}\rightarrow\Omega$ and $\Omega$ is a convex subset of $\mathbb{R}^{d}$. Let $R_{n}$ denote the radius of the space $\Omega_{n}$ which grows at most polynomial order with $n$. If the set of density functions $\mathcal{F}(\Omega_{n})$ is locally Lipschitz in the sense that
$\left|\sqrt{f_{\theta}(\bm{x})}-\sqrt{f_{\vartheta}(\bm{x})}\right|\leq M^{(n)}_{\vartheta,u}(\bm{x})\|\theta-\vartheta\|_{\infty},$
for some function $M^{(n)}_{\vartheta,u}:\mathcal{X}\rightarrow\mathbb{R}$, $\forall\bm{x}\in\mathcal{X}$, $\theta,\vartheta\in\Omega_{n}$, $\|\theta-\vartheta\|_{\infty}\leq u$. Let
$M_{u}^{(n)}:=\sup_{\vartheta\in\Omega_{n}}\|M^{(n)}_{\vartheta,u}\|_{2}$
which is a constant depending on $R_{n}$. If there exists a constant $q\in(0,1)$, such that $\forall u\in(0,\phi_{0})$ for some constant $\phi_{0}$,
$\frac{\log\log M_{u}^{(n)}}{\log n}\leq q,$ (11)
then there exists positive sequence $\phi_{0}>\phi_{n}\rightarrow 0$ with $\phi_{n}>n^{-\tfrac{1-q}{4}}$ such that (10) holds.
Remark 5.
[…]
We will see later in the examples that $M_{u}^{(n)}$ grows with $u$, so usually it suffices to just check (11) for only $u=\phi_{0}$.
Definition 6.
Define for $\alpha\in(0,1]$,
$\rho_{p,\alpha}(\theta,\vartheta):=\frac{1}{\alpha}\int\left[\left(\frac{k_{p}(\bm{x};\theta)}{k_{p}(\bm{x};\vartheta)}\right)^{\alpha}-1\right]k_{p}(\bm{x};\theta)\mathtt{d}\bm{x}.$
For $t>0$, $p\in(0,1)$, define the $\rho_{p,\alpha}$ neighbourhood around $\theta^{*}$ by
$S_{p,\alpha}(t):=\{\theta:\rho_{p,\alpha}(\theta^{*},\theta)\leq t\}.$
Assumption 2.
The positive sequence $\phi_{n}\rightarrow 0$ satisfies satisfies
$\mathbb{P}_{\pi_{0}}(S_{0,\alpha}(t_{n}))\geq c_{3}e^{-2nt_{n}},$ (12)
for some constants $\alpha\in(0,1]$, $c_{3}>0$, where $t_{n}:=c_{1}\phi_{n}^{2}/8$ and $\tfrac{nt_{n}}{\log n}>\max\left\{\tfrac{1}{\alpha},\tfrac{c_{1}}{8c_{2}}\right\}$, $\forall n>0$, $c_{2}$ is a constant in Lemma 9.
Assumption 3.
The positive sequence $\phi_{n}\rightarrow 0$ and $\Omega_{n}\rightarrow\Omega$ satisfy
$\mathbb{P}_{\pi_{0}}(\Omega_{n}^{c})\leq c_{4}\exp(-2c_{1}n\phi_{n}^{2})$ (13)
for some constant $c_{4}>0$.
Assumptions 2 and 3 control the prior concentration on the center and the tails respectively. The first three assumptions are needed in order to control the $\delta$ term in (3). Note that Assumptions 1-3 feature the same symbol $\phi_{n}$ for the sequence. In the proof for the main result, we need the same sequence $\phi_{n}$ to satisfy these three assumptions simultaneously.
Definition 7.
Denote that supremum of the directional derivative of $f(\bm{x};\theta)$ in $\theta$ at $(\bm{x},\theta)$ by ~$\tilde{\nabla}_{\theta}f(\bm{x};\theta)$,
~$\tilde{\nabla}_{\theta}f(\bm{x};\theta)=\sup_{\|\vec{u}\|_{2}=1}\lim_{\xi\downarrow 0}\frac{f(\bm{x};\theta+\xi\vec{u})-f(\bm{x};\theta)}{\xi}.$
Assumption 4.
All the directional derivatives of the density function $f(\bm{x};\theta)$ in $\theta$ exist and are bounded, and the contamination density $g$ has a tail no lighter than ~$\tilde{\nabla}_{\theta}f(\bm{x};\theta)$ in every direction, i.e., for any fixed $\theta\in\Omega$,
$\sup_{\bm{x}\in\mathcal{X}}\left\|\frac{\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{g(\bm{x})}\right\|_{2}<\infty,$
Remark 6.
We assume a weaker condition than total differentiability to accommodate the use of absolute value ||$\cdot$|| ⋅ | in density functions, e.g., the Laplace distribution.
When $g$ is continuous and has full support over the support of ~$\tilde{\nabla}_{\theta}f$, then the assumption holds trivially whenever $g$ has a suitably heavy tail.
Assumption 5.
For every $\psi>0$,
$\inf_{\|\theta-\theta^{*}\|_{2}>\psi}h(f(\cdot;\theta),f(\cdot;\theta^{*}))>0$ (14)
Remark 7.
Assumption 5 is similar to one of the conditions in Thm 9.13 Wasserman (2004), but stronger, since we stated it in Hellinger distance rather than in KL divergence. The condition directly implies the identifiability of the model. See also Condition 2 in the lecture notes Wasserman (2020).
Proposition 4.
If Assumption 5 holds, then for any sequence $\phi_{n}\rightarrow 0$, and decreasing sequence $1>p_{n}\rightarrow 0$, there exists sequence $\psi_{n}\rightarrow 0$ such that
$h(k_{p_{n}}(\cdot;\theta),k_{p_{n}}(\cdot;\theta^{*}))\leq\phi_{n}\implies\|\theta-\theta^{*}\|_{2}\leq\psi_{n}.$ (15)
Theorem 5 ($(\epsilon,\delta)$-Privacy).
Suppose that there exists a decreasing sequence $\phi_{n}>0$ that satisfies the conditions in Assumptions 1-3 and that Assumptions 4, 5 also hold. For $p_{0}\in(0,1)$ fixed, there exists bounded positive constants (or sequence of constants) $\{T_{n}\},c_{1},c_{2},\dots,C_{1},C_{2},\dots<\infty$ such that $\forall S\in\mathcal{B}(\Omega)$ and any sequence $p_{0}\geq p_{n}\downarrow 0$,
$\mathbb{P}_{\theta^{*}}\Bigl{(}\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq e^{\epsilon_{n}}\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}\Bigr{)}\geq\begin{dcases}1-C_{1}e^{-nc_{5}\phi^{2}_{n}},&\text{for }nc_{2}\phi_{n}^{2}\geq\log 2,\\ 1-(C_{2}+C_{3}(\log\phi_{n})^{2})e^{-nc_{5}\phi^{2}_{n}},&\text{otherwise}\end{dcases}$ (16)
with
$\epsilon_{n}:=\varepsilon_{n}+\log(\delta_{n}+\exp(\varepsilon_{n})),\quad\varepsilon_{n}:=2(1-p_{n})T_{n}\psi_{n}/p_{n},\quad\delta_{n}:=\tfrac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}},$
where $\mathbb{P}_{\theta^{*}}$ is the data generating measure for the pair of neighbouring datasets of size $n$, $\bm{X}$ and $\bm{Z}$, with entries generated from the true likelihood $k_{p}(\cdot;\theta^{*})$. In addition, $t_{n}=c_{1}\phi_{n}^{2}/8$, $\psi_{n}$ is defined in Proposition 4, and $c_{5}=\min\{c_{1}\alpha/8,c_{2}\}$. In other words,
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp(\epsilon_{n})\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}$
holds for all large $n$ $\mathbb{P}_{\theta^{*}}$-almost surely.
The takeaway of Theorem 5 is that as $n\rightarrow\infty$, we may choose a sequence of $p_{n}\rightarrow 0$ such that $\epsilon_{n},\delta_{n}\rightarrow 0$ and for $n$ large enough the failure probability falls into the first category of (16), that is $<C_{1}\exp(-nc_{5}\phi_{n}^{2})\rightarrow 0$.
Sketch Proof.
Proposition 2 gives us some clue for how to upperbound the $\epsilon_{n}$ and $\delta_{n}$. Recall that
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\eta_{n}\left[\eta_{n}+\frac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}$
where
$\eta_{n}=\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}.$
Under Assumption 4 and 5, we can prove that $\eta_{n}$ is always finite with $\eta_{n}\rightarrow 1$. We can also characterise its convergence rate (see Lemma 15).
Then we prove the following:
  • In Lemma 13, we prove, under Assumption 2 that
    $\mathbb{P}\left(m_{n}(\Omega,\bm{X})\leq\frac{c_{3}}{2}e^{-4nt_{n}}\right)\leq 2e^{-2n\alpha t_{n}}.$
  • By Lemma 9, Assumption 1 implies that
    $\mathbb{P}^{*}_{\theta^{*}}\left(m_{n}(A_{n}^{c}\cap\Omega_{n},\bm{X})\geq e^{-c_{1}n\phi_{n}^{2}}\right)\leq C^{**}\exp(-c_{2}n\phi_{n}^{2})$
    for some constant $C^{**}$ we specify in the lemma. The proof is provided in the supplementary material.
  • Then Lemma 14 shows that
    $\mathbb{P}\left(m_{n}(A_{n}^{c}\cap\Omega_{n}^{c},\bm{X})\geq e^{-c_{1}n\phi_{n}^{2}}\right)\leq c_{4}e^{-c_{1}n\phi_{n}^{2}},$
    then consequently
    $\mathbb{P}\left(m_{n}(A_{n}^{c},\bm{X})\geq 2e^{-c_{1}n\phi_{n}^{2}}\right)\leq C^{**}e^{-c_{2}n\phi_{n}^{2}}+c_{4}e^{-nc_{1}\phi_{n}^{2}}.$
  • The relevant proofs are contained in the Appendix A, apart from the proof for Lemma 9 in the supplement.
Since we assume that $\bm{Z}$ also follows the same data generating measure, $m_{n}(A_{n}^{c},\bm{Z})$ shares the same upper bound as $m_{n}(A_{n}^{c},\bm{X})$. Now, every term in the inequality has a probabilistic upper bound, combining these bounds and the theorem is proved. ∎
Lemmas 10, 12, Proposition 4 are intermediate results such that whenever the Assumptions 1 and 2 hold with respect to uncontaminated likelihood $f(\cdot;\theta)$, the same also hold for the actual contaminated likelihood $k_{p}(\cdot;\theta)$, and in practice, one only need to check the assumptions for the case $p=0$.
Theorem 6 (Fisher Information).
Let $p_{n}\rightarrow 0$ be a sequence of positive real numbers. Let $I_{p}(\theta)$ denote the Fisher information matrix for a single data point with respect to the density $k_{p}(\cdot;\theta)$ and $I(\theta)$ denote the Fisher information matrix with respect to $f(\cdot;\theta)$. Then
$I_{p_{n}}(\theta^{*})\rightarrow I(\theta^{*}).$
Proof.
The proof is in Appendix B in the supplement. ∎
Proof of Theorem 1.
Under Assumptions 1-5, Theorem 5 holds for any decreasing $p_{n}\rightarrow 0$, $p_{n}\leq p_{0}<1$, especially for $n$ large enough,
$\mathbb{P}\Bigl{(}\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp(\epsilon_{n})\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}\Bigr{)}\geq 1-Ce^{-nc_{5}\phi_{n}^{2}},$
where
$\epsilon_{n}:=\varepsilon_{n}+\log(\delta_{n}+\exp(\varepsilon_{n})),\quad\varepsilon_{n}:=2(1-p_{n})T_{n}\psi_{n}/p_{n},\quad\delta_{n}:=\tfrac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}},$
Thus by picking $p_{n}\rightarrow 0$ such that $\psi_{n}/p_{n}\rightarrow 0$ as $n\rightarrow\infty$, both $\delta_{n}$ and $\epsilon_{n}$ converges to zero. Thus the algorithm is also asymptotically differentially private for any pair of $(\epsilon,\delta)$. Since $p_{n}\rightarrow 0$, the convergence in the Fisher information follows from Theorem 6. ∎
Our results address the case where samples are drawn exactly from the posterior distribution. When the algorithm of interest is a Markov Chain Monte Carlo algorithm (Robert, 1999), i.e., does not output i.i.d. samples, our result indicates the level of differential privacy of such algorithm when it reaches the stationary state. It is important to note that when the Markov chain has converged, the algorithm is irrelevant, and the level of differential privacy only depends on the target distribution.
3.3 Remarks on Assumptions
Remark 8.
The Assumptions 1-3 resembles the set of assumptions in Theorem 8.9 of Ghosal and Van der Vaart (2017). In particular, our Assumption 3 is equivalent to their Assumption (iii). our Assumption 1 is similar to their Assumption (ii) with the semi-metric $d$ replaced by Hellinger distance $h$ and conditioning on a bracketing number instead of covering number. Finally, our Assumption 2 is similar to their Assumption (i) in the sense that both conditions on the prior mass around a shrinking neighbourhood around the true parameter $\theta^{*}$ are sufficiently large.
We adopt the current set of assumptions in order to use the proof techniques of Shen and Wong (1994); Wong and Shen (1995); Shen and Wasserman (2001), which characterizes the rate of convergence the quantity $\mathbb{P}_{n}(A_{n}^{c})$ with a known failing rate, i.e., (16) type bound, which is not available by the Theorems of Ghosal and Van der Vaart (2017).
Remark 9.
One sufficient condition for Assumption 1 to hold is the following:
$H_{[]}(u,\mathcal{F}(\Omega_{n}))\leq c^{2}n\phi_{n}^{2},\forall u\in(\phi_{n}^{2}/2^{10},\phi_{n}).$
By Lemma 3, we know the above will hold for most likelihood functions $f(\cdot;\theta)$ when $n$ is large enough. However, due to the constant $c$ being quite small in magnitude, see Lemma B.2 in the supplement, one usually need $n$ to be unrealistically large for Theorem 5 to hold. Thus despite having derived its convergence rate in a non-asymptotic fashion, the result is better deemed as an asymptotic one. We will show in the next section with empirical simulation results that Theorem 5 does give a good intuition as to how $\epsilon_{n}$ and $\delta_{n}$ converge.
Both Assumptions 2 and 3 are conditions on the prior distribution. Let $\bar{S}(r):=\{\theta\in\Omega:\|\theta^{*}-\theta\|_{2}\leq r\}$ denote the $L_{2}$-ball of radius $r$ around $\theta^{*}$. It is possible to show that under mild conditions, fixing a constant $c_{r}$, the set $S_{0,\alpha}(t_{n})\cap\bar{S}(c_{r})$ contains a $L_{2}$-ball with radius proportional to $t_{n}$.
Lemma 7.
Let $S_{0,\alpha}(t_{n}):=\{\theta:\rho_{0,\alpha}(\theta^{*},\theta)\leq t_{n}\}$ defined as in Definition 2. Suppose that there exists $\alpha\in(0,1]$, and $c_{r}<\infty$ such that
$\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\int\left[\frac{f(\bm{x};\theta^{*})}{f(\bm{x};\theta)}\right]^{\alpha+1}\|\tilde{\bm{\nabla}}_{\theta}f(\bm{x};\tilde{\theta})\|_{2}\mathtt{d}x<M_{c_{r}}$
where $M_{c_{r}}$ is a constant that depends on $c_{r}$, and $\tilde{\bm{\nabla}}$ is supremum directional derivative defined in Definition 7. Then $\bar{S}(t_{n}/M_{c_{r}})\subseteq S_{0,\alpha}(t_{n})$.
Remark 10.
The condition holds for most likelihood functions, since ~$\|\tilde{\bm{\nabla}}_{\theta}f(x;\theta)\|$ is usually integrable already. If $\theta$ does not control the dispersion, then $f(x;\theta^{*})$ and $f(\bm{x};\theta)$ have matching rates of decay and hence the expression is integrable. Otherwise, one usually can choose $c_{r}$ small enough such that the expression remains integrable.
Remark 11.
One may use a similar argument to show that the Hellinger distance and the $L_{2}$ distance are locally related (See Appendix C) so that $\phi_{n}$ and $\psi_{n}$ can be chosen to decay at the same rate. Then, the rate of $p_{n}$ can be chosen freely as long as $\psi_{n}/p_{n}\rightarrow 0$.
Note that we may choose $\Omega_{n}$ in Assumption 3 as an expanding $L_{2}$-ball. Then both Assumptions 2, 3 reduce to computing the prior mass on the $L_{2}$-ball $\bar{S}(r)$, i.e., $\mathbb{P}_{\pi_{0}}(\bar{S}(r))$ for some (sequence of) $r$. In general, Assumption 2 should hold whenever there is sufficient mass around the true parameter $\theta^{*}$, i.e., the prior’s support covers $\theta^{*}$ and Assumption 3 should hold whenever the prior has an exponentially decaying tail.
When the prior has a subexponential tail, Assumption 3 requires the set $\Omega_{n}$ to expand at least exponentially fast with $n$. For instance, it is easy to check that the horseshoe prior satisfies Assumption 3 easily when $\Omega_{n}$ expands with radius $e^{n}$. Following the proof of Lemma 3, one may derive a similar result for $\Omega_{n}$ with radius growing at most $O(e^{n})$ fast, which still hold asymptotically for a wide range of common likelihood functions, e.g., for Gaussian.
The following Lemma is useful when the prior is a standard Gaussian, which also adapts trivially to a spike and slab prior with a Gaussian slab.
Lemma 8.
Let $\pi_{0}(\cdot)$ and $\bar{S}(r)$ be defined as above, we have two separate lower bounds for the prior mass
$\mathbb{P}_{\pi_{0}}(\bar{S}(r))\geq\begin{dcases}\frac{\exp(-\lambda/2)}{\Gamma(d/2)}\gamma(d/2,r^{2}/2)&\text{ for }r\text{ small}\\ \frac{1}{\Gamma(d/2)}\gamma(d/2,R^{2}/2)&\text{ for }r>\|\theta^{*}\|_{2}.\end{dcases}$
where $\lambda=\|\theta^{*}\|_{2}^{2}$, $R=r-\|\theta^{*}\|_{2}$ and
$\gamma(\alpha,r)=\int_{0}^{r}u^{\alpha-1}\exp(-u)\mathtt{d}u$
is the incomplete gamma function.
Remark 12.
For $r$ small, we use the first inequality to verify Assumption 2
$\frac{\gamma(\alpha,r)}{\Gamma(\alpha)}=\,\,$ $\frac{1}{\Gamma(\alpha)}\int_{0}^{r}w^{\alpha-1}e^{-w}\mathtt{d}w$
[…]
where the last step is due to Stirling’s approximation, (see, for instance, (3.9) page 24 of Artin (1964)). Thus $\mathbb{P}_{\pi_{0}}(\bar{S}(r))$ shrinks polynomially with $r$ which is always slower than the exponential shrinkage in Assumption 2 for $n$ large enough.
Remark 13.
For $r$ large, to verify Assumption 3, we use the second inequality with the approximation $\Gamma(\alpha)-\gamma(\alpha,{r})\sim r^{\alpha-1}e^{-r}$, or
$\frac{\gamma(\alpha,r)}{\Gamma(\alpha)}\sim 1-\frac{1}{\Gamma(\alpha)}r^{\alpha-1}e^{-r}$
and hence,
$1-\mathbb{P}_{\pi_{0}}(\bar{S}(r))\preceq r^{d-2}\exp(-(r-\|\theta^{*}\|_{2})^{2}/2).$
where $\sim$,$\preceq$ , ⪯ are
  1. 1.
    $f(x)\sim u(x)\text{ if }f/u\rightarrow 1$ as $x\rightarrow L,$ or the leading term of $f(x)$ is $u(x)$ as $x\rightarrow L$;
  2. 2.
    $f_{n}(x_{n})\preceq u_{n}(x_{n})$ if there exists constant $C>0$ independent of $x_{n}$ and $n$ such that $f_{n}(x_{n})\leq Cu_{n}(x_{n})$.
4 Empirical Assessment of Differential Privacy
It is nontrivial to directly link the $(\epsilon,\delta)$ of the model with the amount of contamination being applied. In other words, it can be hard for the practitioners to directly derive the amount of contamination needed to achieve a certain degree of $(\epsilon,\delta)$ by consulting Theorem 5 along. Thus, this section is devoted to provide guidance on how to empirically estimate $(\epsilon,\delta)$ of the model in practice. Hence, the choice of contamination may be tuned accordingly after the empirical estimation, e.g., increase contamination rate $p_{n}$ if $\epsilon$ is not low enough.
We begin by introducing our pipeline of estimating DP, then followed by some empirical simulations for a few settings of common regression models.
4.1 Estimation Procedure
Estimating $\epsilon$ and $\delta$ mostly follows from the result of Proposition 2. Recall (9) from Proposition 2,
$\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{X})\leq\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{m_{n}(\Omega,\bm{Z})}{m_{n}(\Omega,\bm{X})}\mathbb{P}_{\pi_{n}}(S\cap A_{n}|\bm{Z}),$
where
$d(\bm{x};\theta):=\frac{k_{p_{n}}(\bm{x};\theta)}{k_{p_{n}}(\bm{x};\theta^{*})}=\frac{(1-p_{n})f(\bm{x};\theta)+p_{n}g(\bm{x})}{(1-p_{n})f(\bm{x};\theta^{*})+p_{n}g(\bm{x})}$
Thus
$\exp(\hat{\epsilon}_{n})\leq\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{m_{n}(\Omega,\bm{Z})}{m_{n}(\Omega,\bm{X})}=\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\frac{\mathbb{E}_{\pi_{n-1}}[d(\bm{z};\theta)]}{\mathbb{E}_{\pi_{n-1}}[d(\bm{x};\theta)]}$
where $\pi_{n-1}$ is the posterior distribution on the common $n-1$ data points shared between $\bm{X}$ and $\bm{Z}$. The first fraction can be dealt with separately for $\bm{x}$ and $\bm{z}$, by solving
$\sup_{\bm{x}\in\mathcal{X}}\sup_{\theta\in A_{n}}d(\bm{x};\theta)\text{ and }\inf_{\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}d(\bm{z};\theta).$ (17)
Since $\pi_{n-1}$ is usually only available up to a constant, we consider solving the Monte Carlo approximated optimization problems
$\sup_{\bm{x}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i}),\quad\inf_{\bm{z}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i})$ (18)
where $\theta_{i}$ are drawn from a proposal distribution, e.g., the Laplace approximation of the posterior distribution (see Section 4.6.8.2, pg. 152-153 of Murphy (2022)),
$\theta_{i}\sim\mathcal{N}(\theta_{\text{MAP}},\bm{H}^{-1})$ (19)
where $\theta_{\text{MAP}}$ is the map estimator of the posterior distribution $\pi_{n-1}$ and $\bm{H}$ is the hessian of $-\log(\pi_{n-1}(\theta|\mathcal{D}))$.
For $\delta$, we will estimate $\mathbb{P}_{\pi_{n}}(A_{n}^{c}|\bm{X})$ as an upperbound for $\delta_{n}$. Again, an analytic solution is usually unavailable, so a numerical scheme is required. Here we recommend using the Laplace approximation again to implement an importance sampling scheme by sampling $\theta_{i}\sim\mathcal{N}(\theta_{\text{MAP}},\bm{H}^{-1})$ and compute
$\hat{\delta}_{n}:=\sum_{\theta_{i}\in A_{n}^{c}}w_{i}.$ (20)
There are two user-defined sequences $\phi_{n}$ and $p_{n}$ that define the radius of the neighbourhood $A_{n}$ and the contamination ratio respectively. In the simulations, we choose
$\phi_{n}:=C_{\phi}^{(1)}\exp(-n^{1/C_{\phi}^{(2)}}),\quad p_{n}:=\exp(-n^{1/C_{p}})$ (21)
where $C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p}$ are constants you may find in the relevant figures/descriptions. Finally, we construct the set $A_{n}$ to be a neighbourhood around $\theta^{*}$, e.g., $L_{\infty}$-ball around $\theta^{*}$. Now, $\epsilon_{n}$ and $\delta_{n}$ may be estimated by Algorithm 1.
Remark 14.
There are a few places where the use of $\theta^{*}$ is preferred. For the simulations, we can use $\theta^{*}$ directly, but when $\theta^{*}$ is unknown, an estimation $\hat{\theta}^{*}$ would suffice.
Remark 15.
In order for Assumption 4 to be satisfied for regression settings, one needs to impose a constraint on the size of covariates. In theory, we may allow the size of $\bm{W}$ to grow with $n$, for simplicity, we assume in all the simulations that $\|\bm{W}\|_{\infty}\leq 1$.
input : Constants $C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p}$; Likelihood function $f$; Contamination density $h$; prior $\pi_{0}$; Data set size $n$; Number of repeats $K$; Number of particles $m$; Estimated true parameter $\hat{\theta}^{*}$;
1 Compute $\phi_{n},p_{n}$ in (21);
2 Set $A_{n}:=B_{\phi_{n}}(\hat{\theta}^{*},\infty)$ to be the $L_{\infty}$-ball around $\hat{\theta}^{*}$, i.e., a hypercube of width $2\phi_{n}$.;
3 Solve
$\eta_{n}:=\sup_{\bm{x},\bm{z}\in\mathcal{X}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}$
by solving (17) with respect to $A_{n},p_{n}$;
4 for $k=1,\dots,K$ do
5       if No dataset is present then
6             Generate covariates $\bm{W}:=\{\bm{w}_{1},\dots,\bm{w}_{n}\}$ from a desired distribution, e.g., standard multivariate normal.;
7             Generate true observations $\bm{X}:=\{\bm{x}_{1},\dots,\bm{x}_{n}\}$ according to $\bm{x}_{i}\sim f(\cdot;\theta^{*}|\bm{w}_{i})$;
8
9       Contaminate the observations $\bm{X}$, denoted $\tilde{\bm{X}}$, by replacing each $\bm{x}_{i}$ with probability $p_{n}$ with a draw from $g(\cdot)$;
10       Compute the Laplace approximation of the contaminated posterior distribution $\pi_{n}(\cdot;\tilde{\bm{X}})$ according to (19) ;
11       Sample $\theta_{1},\dots,\theta_{m}$ from the Laplace approximation;
12       Compute normalized importance weights $w_{1},\dots,w_{m}$ with respect to the $\pi_{n}(\cdot;\tilde{\bm{X}})$;
13       Compute $\hat{\delta}_{n}^{(k)}$ in (20);
14       Set $\mathcal{D}$ to be $\tilde{\bm{X}}$ with the last element discarded;
15       Compute the normalized weight $\tilde{w}_{i}$ with respect to $\pi_{n-1}(\cdot;\mathcal{D})$;
16       Solve for
$\alpha:=\sup_{\bm{x}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i}),\quad\beta:=\inf_{\bm{z}}\sum_{i=1}^{m}w_{i}d(\bm{x};\theta_{i})$
using $\theta_{1},\dots,\theta_{m}$ and $\tilde{w}_{1},\dots,\tilde{w}_{m}$;
17       Set $\hat{\epsilon}_{n}^{(k)}:=\eta_{n}+\log(\alpha)-\log(\beta)$;
18
19 end for
output : Estimated $\hat{\epsilon}_{n}^{(k)}$, $\hat{\delta}_{n}^{(k)}$, $k=1,\dots,K$
Algorithm 1 An empirical procedure to estimate $\epsilon$ and $\delta$ for a given set of parameters $A_{n},p_{n}$.
4.2 Simulation Setups and Results
Let $\bm{X}=\{X_{1},\dots,X_{n}\}\in\mathbb{R}^{n}$ be the set of observations and $\bm{W}\in[-1,1]^{n\times d}$ be the covariate matrix associated with $\bm{X}$, i.e., the $i$-th row of $\bm{W}$ is associated with the $i$-th observation $X_{i}$. By default, we assume the first column of $\bm{W}$ is a vector of $1$. Let $\theta\in\mathbb{R}^{d}$ denote the parameter and the model likelihood is given by a density function $f(X_{i};\theta,\bm{w}_{i})=f(X_{i},\theta^{\top}\bm{w}_{i})$, such that $\int_{\mathbb{R}}f(x;\theta,\bm{w})\mathtt{d}x=1$ for any pair of $\theta,\bm{w}\in\mathbb{R}^{d}$. Let $\theta^{*}$ denote the true parameter value, which in practice may be replaced by an estimation $\hat{\theta}$, e.g., using the maximum a posteriori estimator.
For the cases where contamination is required, we choose a contamination density $g(x;\bm{w})$ and replace each $X_{i}$ with probability $p_{n}$ by a random sample from $g(\cdot;\bm{w}_{i})$. Note that the contamination density is allowed to depend on the covariate $\bm{w}$, see the example below for the linear regression case.
When contamination is present, the Bayesian posterior will be computed with respect to the contaminated likelihood
$k_{p_{n}}(x;\theta,\bm{w})=(1-p_{n})f(x;\theta,\bm{w})+p_{n}\,g(x;\bm{w}).$
The prior $\pi_{0}$ on $\theta$ is always set as a zero mean Gaussian prior with the appropriate dimension. The covariance is a diagonal matrix with all diagonal terms equal to $100$.
The above estimation procedure is applied to the following three models:
  1. 1.
    Bayesian Linear Regression: ($C_{\phi}^{(1)}=5$)
    The regression model is given by
    $X_{i}=\theta^{*\top}\bm{w}_{i}+e_{i}$
    where $e_{i}\sim\mathcal{N}(0,1)$. The likelihood function $f(X;\theta,\bm{w})$ is given by
    $f(X;\theta,\bm{w})=\frac{1}{2\pi\sigma^{2}}\exp\left(-\frac{(X-\theta^{\top}\bm{w})^{2}}{2}\right).$
    The contamination density is chosen as $g(X_{i};\bm{w}_{i})\sim T_{5}(\theta^{*\top}\bm{w}_{i})$, a shifted Student’s t-distribution with degrees of freedom $5$, given by the following expression
    $g(x;\bm{w})=\frac{\Gamma(3)}{\sqrt{3\pi}\Gamma(\tfrac{5}{2})}\left(1+\frac{(x-\theta^{*\top}\bm{w})^{2}}{5}\right)^{-3}$
    where $\sigma$ is the standard deviation of the model noise for convenience. In this setup, a heavy-tailed contamination is required in order to remove the compactness constraint on the observation space.
  2. 2.
    Bayesian Logistic Regression: ($C_{\phi}^{(1)}=10$)
    The regression model is given by $X_{i}\sim\text{Ber}\left((1+e^{-\theta^{*\top}\bm{w}_{i}})^{-1}\right)$ where $X\in\{0,1\}$ with likelihood function
    $f(X;\theta,\bm{w})=\left(1+e^{-\theta^{*\top}\bm{w}_{i}}\right)^{-X}\left(1+e^{\theta^{*\top}\bm{w}_{i}}\right)^{-1+X}.$
    Since the observation space is compact, $\epsilon$ is bounded in this setup. However, as we will see in Section 16, contamination should still be adopted. To contaminate the observations in this case, we replace each observation $X_{i}$ with probability $p_{n}$ by a draw from $\text{Ber}(\tfrac{1}{2})$. Thus the contaminated density is $k_{p_{n}}(X;\theta,\bm{w})=(1-p_{n})f(X;\theta,\bm{w})+0.5p_{n}.$
  3. 3.
    Bayesian Cauchy Regression: ($C_{\phi}^{(1)}=8$)
    The regression model is given by $X_{i}=\theta^{*\top}\bm{w}_{i}+e_{i}$ where $e_{i}\sim\text{Cauchy}(0)$ with likelihood function
    $f(X;\theta,\bm{w})=\frac{1}{\pi\left[1+\left(X-\theta^{\top}\bm{w}\right)^{2}\right]}.$
    The simulations are done without contamination as $d(X;\theta)$ has bounded sensitivity.
Each of the above setups is simulated with parameter dimension $=51$ and the estimated $(\epsilon_{n},\delta_{n})$ results are plotted in Figure 1. The three rows each represent the results for Linear regression, Logistic regression (with contamination) and Cauchy regression respectively.
For each combination of $(C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p})$, the simulation is repeated for $48$ times. For each combination of $(C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p})$, the average $\hat{\epsilon}_{n},\hat{\delta}_{n}$ over all repetitions, and the maximum $\hat{\epsilon}_{n},\hat{\delta}_{n}$ over all repetitions are plotted. The max curves and average curves are distinguished by the point shape, circle for average and triangle for max. The curves are plotted against $n$, the number of data points in the dataset. We see in all cases, the max curve and the average curve shrink at a similar rate.
Figure 1: Plots of estimated $(\epsilon_{n},\delta_{n})$ for three different model setups, with parameter dimension $51$. The average and maximum across all repeats are plotted and distinguished by dot shape. Each colour and line type pair corresponds to the result of one choice of the constant.
(a) Linear regression case
(b) Logistic regression case
(c) Cauchy regression case
(d) Uncontaminated logistic regression
Remark 16 (Contaminated Logistic Regression).
Similar to the Cauchy case, Logistic regression is differentially private without any contamination. In Figure 1, the $\hat{\epsilon}_{n}$ and $\hat{\delta}_{n}$ are plotted for the contaminated case and uncontaminated case in the 2nd and the 4th row respectively. Again, the max and average curves are computed for each combination of the constants $(C_{\phi}^{(1)},C_{\phi}^{(2)},C_{p})$ across multiple repetitions. We see a significant improvement through contamination, where $\hat{\epsilon}_{n}$ in the contaminated setup is only $1/10$ of the $\hat{\epsilon}_{n}$ in the uncontaminated setup while $\hat{\delta}_{n}$ curves are similar.
Remark 17 (Effect of Output Dimension).
We also ran the same setups but with parameter dimension $5$ and the results are plotted in Figure 2. We see that with lower dimensions, $\hat{\epsilon}_{n}$ are smaller in all cases. This is inevitable since a higher dimensional sample contains more information about the data set than a lower dimensional one.
Figure 2: Plots of estimated $(\epsilon_{n},\delta_{n})$ for the three setups with parameter dimension $5$.
(a) Linear regression case
(b) Logistic regression case
(c) Cauchy regression case
5 Examples
We have seen in the previous section that empirically the approximated $\hat{\epsilon}_{n},\hat{\delta}_{n}$ follows the decreasing trend as in Theorem 5. In this section, we give a few examples that Theorem 5 is applicable theoretically. The verification of Assumptions 1 to 3 across different models shares a few common procedures. For Assumption 4 to hold for regression models (with covariates), one needs to control the leverage of the covariates $\bm{W}$. It suffices to assume $\|\bm{W}\|_{\infty}\leq n^{\iota}$, for some $\iota\in(0,1)$ that depends on the choice of $\phi_{n}$. However, for simplicity, we will still assume $\|\bm{W}\|_{\infty}\leq 1$, as we did in the simulations in Section 4. With $\|\bm{W}\|_{\infty}\leq 1$, Assumption 4 can be trivially satisfied by correctly choosing a heavy-tailed contamination density $g$, and Assumption 5 is usually easy to check for linear models.
Due to the length, we will only present the Bayesian quantile regression case in this section. We show three other cases in Appendix C that is contained in the supplement (Hu et al., 2024).
5.1 Bayesian Quantile Regression
In quantile regression, one would like to estimate the $\tau$-th quantile, denoted $q_{\tau}(X_{i}|\bm{W}_{i})$ of the random variable $X_{i}$ conditioned on the covariates $\bm{W}_{i}$. Due to Koenker and Bassett Jr (1978), solving the quantile regression problem is equivalent to solving the following optimization problem:
$\min_{\theta}\sum_{i}Q_{\tau}(x_{i}-\theta^{\top}\bm{w}_{i}),\quad Q_{\tau}(u)=\frac{1}{2}(|u|+(2\tau-1)u)$
This is equivalent to maximizing the likelihood
$\max_{\theta}\prod_{i}f_{\tau}(x_{i};\bm{w}_{i},\theta)$
where the probability density function of $X$ given $\tau,\theta,\bm{w}$ is
$f^{\tau}_{\theta}(x|\bm{w})=\tau(1-\tau)\exp\left(-Q_{\tau}(x-\theta^{\top}\bm{w})\right)$
describes the asymmetric Laplace distribution. Please refer to Yu and Moyeed (2001) for more details on Bayesian quantile regression. Firstly, we check Assumption 1 by using Proposition 3. We showed that, by Lemma 27, for any $\theta,\vartheta\in\Omega_{n}:=\{\theta\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n}\}$, such that $\|\theta-\vartheta\|_{\infty}\leq r$,
$\left|\sqrt{f^{\tau}_{\theta}(x|\bm{w})}-\sqrt{f^{\tau}_{\vartheta}(x|\bm{w})}\right|\leq M_{\vartheta,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty}$
where $\tau^{*}=\max\{\tau,1-\tau\}$, $M_{n}=R_{n}+\|\theta^{*}\|_{2}$ and
$M_{\vartheta,r}(x,\bm{w})=\frac{d}{2}\exp\left(-\frac{1-\tau^{*}}{2}\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|\right)\|\bm{w}\|_{2}.$
Then by Lemma 28, assuming $\|\bm{w}\|_{\infty}\leq 1$,
$\|M_{\vartheta,r}\|_{2}^{2}=\frac{d^{2}}{2(1-\tau^{*})}\|\bm{w}\|_{2}^{2}+\frac{d^{2}r}{2}\|\bm{w}\|_{2}^{3}=O(1).$
Thus Assumption 1 can be satisfied with the choice $\phi_{n}=n^{-1/q},R_{n}=n^{k}$ for some $q>4$ and $k>0$ to be chosen for Assumption 3. It is not hard to check that the likelihood function satisfies the condition in Lemma 7, when $\theta$ is close enough to $\theta^{*}$, i.e., within $L_{2}$-distance of some constant $c_{r}$. Thus $S_{0,\alpha}(t_{n})$ contains an $L_{2}$-neighbourhood around $\theta^{*}$ with its radius proportional to $t_{n}$ up to a constant depending on $\theta^{*}$ and $c_{r}$. Then Assumption 2 is satisfied by applying Lemma 8 and Remark 12. For Assumption 3, we consider for $R_{n}$ large and apply Remark 13,
$\mathbb{P}_{\pi_{0}}(\Omega_{n}^{c})=1-\mathbb{P}_{\pi_{0}}(\Omega_{n})\sim R_{n}^{d-2}e^{-R_{n}^{2}/2}.$
and we need
$R_{n}^{2}=n^{2k}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q}$
which can be satisfied by choosing $k>(q-2)/2q$, for instance, $q=4$, $k>\tfrac{1}{4}$.
Finally, Lemma 29 also shows that $\phi_{n}$ and $\psi_{n}$ in Proposition 4 can be chosen to diminish at the same rate, similar to the conclusions as in the previous examples. Thus, we can choose $\psi_{n}\propto n^{-1/q}$ and $p_{n}=n^{-1/2q}$ which gives
$\epsilon_{n}\sim n^{-1/2q},\qquad\delta_{n}\sim\exp(-\tfrac{1}{2}c_{1}n^{1-2/q})$
where $c_{1}$ is a fixed constant and $q>4$.
6 Discussion
Differential privacy mechanisms in statistical inference can be roughly divided into two categories: 1. Local models, where data are privatized before collection and inference is conducted on the perturbed dataset; 2. Central models, where data are collected and processed without noise but the result is privatized before release.
The focus of the ”local model”, or ”local differential privacy” as defined in Erlingsson, Pihur and Korolova (2014), quantifies the level of differential privacy provided by the mechanism on ”datasets” with only one data point. Thus contamination can be done individually before the dataset is gathered, Huber’s contamination model is a popular approach to local differential privacy. Our work, however, considers differential privacy from a central model perspective despite using a local privatization mechanism. Notably, the local differential privacy level only depends on the contamination probability $p_{n}$, but $p_{n}\rightarrow 0$ as $n\rightarrow\infty$. From the local model perspective, our setup will be asymptotically non-private as $n\rightarrow\infty$. However, from the central model perspective, with a trusted curator, our approach can ensure full differential privacy with asymptotically zero contamination rate.
Huber’s contamination model for robust estimation by contaminating the dataset with a heavy-tailed distribution also introduces differential privacy to posterior sampling similar to central models that directly inject Laplace/Gaussian noise into the output. By contaminating the dataset with a heavy-tailed distribution, we can lift the assumption on bounded observation space. We further extend the result, through a probabilistic argument, to lift the bound on the parameter space and upper-bound the rate of convergence for $\epsilon$ and $\delta$.
Arguably, privacy through contamination is more robust since the parameter/observation spaces are allowed to be unbounded and the introduced noise is mostly unaffected by the change in the size of the space, as shown in the simulations. Through simulation, a practitioner can estimate the level of differential privacy without having to hand-compute all the constants in the theorem. Despite having a couple of non-trivial assumptions for the theorem, we show in Section 5, how the assumptions could be verified mathematically and how the convergence rate of $\epsilon$ and $\delta$ can be derived under a few common regression setups.
Our main result (Theorem 5) in principle holds uniformly for all data size $n$ as long as the assumptions are satisfied. However, one should perceive it as an asymptotic result since $n$ has to be large for some assumption to hold in general. Shall the constants be improved will extend the practicality of our results. In addition, our main result is built primarily on the contraction of the posterior distribution. It is plausible to extend our result to other interesting models like misspecified models, infinite-dimensional models, etc, when the compatible posterior contraction results are present.
The underlying assumption of our analysis is that i.i.d. samples are drawn exactly from the target distribution. Applying results from Minami et al. (2016) can extend the our bounds non-exact samplers if the algorithm has an established convergence bound. The analysis of Markov chain convergence is an interesting area of research on its own with extensive literature, e.g., Andrieu et al. (2022). Thus we would like to leave the relevant analysis as future work and assume a ”good” sampler is readily available.
Appendix A Proofs for Section 3.2
Lemma 9 (Theorem 1 (Wong and Shen, 1995)).
Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be i.i.d. random variables following density $k_{p}(\cdot;\theta^{*})$. Let $d(\bm{x},\theta)$ denote the likelihood ratio between $\theta$ and $\theta^{*}$ at $\bm{x}$, $d(\bm{x},\theta)=k_{p}(\bm{x};\theta)/k_{p}(\bm{x};\theta^{*})$. Then $\forall\phi_{n}>0$, $\Omega_{n}\subset\Omega$, if
$\int_{\phi_{n}^{2}/2^{10}}^{\phi_{n}}H_{[]}^{1/2}\left(u,\mathcal{K}_{p}(\Omega_{n})\right)\mathtt{d}u\leq c\sqrt{n}\phi_{n}^{2},$ (22)
then
$\mathbb{P}^{*}_{\theta^{*}}\left(\sup_{\theta\in A_{n}^{c}\cap\Omega_{n}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq\exp({-c_{1}n\phi_{n}^{2}})\right)\leq(1+L(\phi_{n}))(1+N(\phi_{n}))\exp({-c_{2}n\phi_{n}^{2}})$
where $\mathbb{P}^{*}_{\theta^{*}}$ denotes the outer measure with respect to $\bm{X}_{i}\sim k_{p}(\cdot;\theta^{*})$, $A_{n}:=\{\theta\in\Omega:h(\theta^{*},\theta)\leq\phi_{n}\}$, and $c,c_{1},c_{2}$ are positive constants independent of $n$ and $\phi_{n}$.
If in addition that $\exp\left(-c_{2}n\phi_{n}^{2}\right)\leq\frac{1}{2},$ then
$\mathbb{P}^{*}_{\theta^{*}}\left(\sup_{\theta\in A_{n}^{c}\cap\Omega_{n}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq\exp({-c_{1}n\phi_{n}^{2}})\right)\leq 4\exp({-c_{2}n\phi_{n}^{2}}).$
Proof.
See Appendix B.4 in the supplementary material. In our choice of constants, we have $(1+L(\phi_{n}))(1+N(\phi_{n}))\leq 20\max\{-\log_{2}(\phi_{n}),(\log_{2}(\phi_{n}))^{2}\}$. ∎
Lemma 10.
Let $p\in[0,1)$, $A\subset\Omega$, $\mathcal{F}(A):=\{f(\cdot;\theta):\theta\in A\}$ and $\mathcal{K}_{p}(A):=\{k_{p}(\cdot;\theta);\theta\in A\}$. Suppose that $1>p\geq p^{\prime}$, then
$H_{[]}\left(u\sqrt{\frac{1-p}{1-p^{\prime}}},\mathcal{K}_{p}(A)\right)\leq H_{[]}(u,\mathcal{K}_{p^{\prime}}(A)).$ (23)
Setting $p^{\prime}=0$, and we have
$H_{[]}\left(u,\mathcal{K}_{p}(A)\right)\leq H_{[]}\left(u\sqrt{1-p},\mathcal{K}_{p}(A)\right)\leq H_{[]}(u,\mathcal{F}(A)).$ (24)
Lemma 11 (Shen and Wasserman (2001)).
Let $t_{n}$ be a sequence of positive numbers and $\alpha\in(0,1]$. For any fixed $p\in[0,1)$,
$\mathbb{P}\left(m_{n}(\Omega,\bm{X})\leq\frac{1}{2}\mathbb{P}_{\pi_{0}}(S_{p,\alpha}(t_{n}))e^{-2nt_{n}}\right)\leq 2e^{-n\alpha t_{n}}.$
Proof.
See Lemma 2 of Shen and Wasserman (2001), the proof holds for any fixed $p\in[0,1)$. ∎
Lemma 12.
For $\alpha\in(0,1]$, $S_{p,\alpha}(t)=\{\rho_{p,\alpha}(\theta^{*},\theta)\leq t)\}$, then $S_{0,\alpha}(t)\subseteq S_{p,\alpha}(t)$.
Lemma 13.
Under Assumption 2, we have for any $p\in[0,1)$
$\mathbb{P}\left(m_{n}(\Omega,\bm{X})\leq\frac{c_{3}}{2}e^{-4nt_{n}}\right)\leq 2e^{-n\alpha t_{n}}.$
Proof.
By Lemma 12 and Assumption 2,
$\mathbb{P}_{\pi_{0}}\Bigl{(}S_{p,\alpha}(t_{n})\Bigr{)}\geq\mathbb{P}_{\pi_{0}}\Bigl{(}S_{0,\alpha}(t_{n})\Bigr{)}\geq c_{3}e^{-2nt_{n}}.$
Then apply Lemma 11 with the above inequality to give the bound. ∎
Lemma 14.
Recall that $A_{n}:=\{\theta\in\Omega:h(\theta,\theta^{*})\leq\phi_{n}\}$. Under Assumptions 1 and 3, we have for any $p\in(0,1)$, $n>0$,
$\mathbb{P}\left(m_{n}(A_{n}^{c},\bm{X})\geq 2e^{-c_{1}n\phi_{n}^{2}}\right)\leq C(n,\phi_{n})e^{-c_{2}n\phi_{n}^{2}}+c_{4}e^{-nc_{1}\phi_{n}^{2}}.$
where
$C(n,\phi_{n})=\begin{dcases}4,&\text{for }nc_{2}\phi_{n}^{2}\geq\log 2.\\ 20\max\{-\log_{2}(\phi_{n}),(\log_{2}(\phi_{n}))^{2}\},&\text{otherwise}\end{dcases}$
Lemma 15.
Under Assumptions 4 and 5, there exists a sequence $T_{n}<\infty$ such that for any sequence $p_{n}\rightarrow 0$
$\sup_{x\in\mathbb{R}^{d}}\sup_{\theta\in A_{n}}|\log d(\bm{x};\theta)|\leq\frac{1-p_{n}}{p_{n}}T_{n}\psi_{n}.$
where $d(\bm{x};\theta)=k_{p_{n}}(\bm{x};\theta)/k_{p_{n}}(\bm{x};\theta^{*})$. Thus in (7), $\eta_{n}=\exp({2(1-p_{n})T_{n}\psi_{n}/p_{n}})$.
Proof for Theorem 5.
By Lemma 13 and 14, and recall $t_{n}=c_{1}\phi_{n}^{2}/8$ defined in Assumption 2,
$\mathbb{P}\left(\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}\geq\frac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}}\right)\leq$ $\,\mathbb{P}\left(\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}\geq\frac{4}{c_{3}}e^{-c_{1}n\phi_{n}^{2}+4nt_{n}}\right)$
$\leq$  $\,2e^{-nc_{1}\alpha\phi_{n}^{2}/8}+C(n,\phi_{n})e^{-c_{2}n\phi_{n}^{2}}+c_{4}e^{-c_{1}n\phi_{n}^{2}}$
Recall from (7) that for any measurable $S\subset\Omega$
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\eta_{n}\left[\eta_{n}+\frac{m_{n}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})},$
and from Lemma 15 that $\eta_{n}\leq e^{2(1-p_{n})T_{n}\psi_{n}/p_{n}}.$ Let $\delta_{n}=\tfrac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}}$, $\varepsilon_{n}=2(1-p_{n})T_{n}\psi_{n}/p_{n}$ and $\epsilon_{n}=\varepsilon_{n}+\log(\delta_{n}+\exp(\varepsilon)).$ Note that since $\bm{X}$ and $\bm{Z}$ has the same cardinality, we apply Lemma 14 again to $m_{n}(A_{n}^{c},\bm{Z})$, we have
$\mathbb{P}\Bigl{(}\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq e^{\epsilon_{n}}\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}\Bigr{)}\geq$  $\,1-\mathbb{P}\left(\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n}(\Omega,\bm{X})}\geq\frac{4}{c_{3}}e^{-\frac{1}{2}c_{1}n\phi_{n}^{2}}\right)$
$\geq$  $\,1-2e^{-nc_{1}\alpha\phi_{n}^{2}/8}-C(n,\phi_{n})e^{-c_{2}n\phi_{n}^{2}}-c_{4}e^{-c_{1}n\phi_{n}^{2}}$
$\geq$  $\,1-C^{\prime}(n,\phi_{n})e^{-nc_{5}\phi_{n}^{2}}$
where $c_{5}=\min\{c_{1}\alpha/8,c_{2}\}$ and $C^{\prime}(n,\phi_{n})=2+C(n,\phi_{n})+c_{4}$. Note by definition of $C(n,\phi_{n})$ from Lemma 14, $C^{\prime}(n,\phi_{n})=20(\log_{2}n)^{2}$. On the other hand, if $n$ is large enough, then $C^{\prime}(n,\phi_{n})=6+c_{4}$ is independent of $n$, and the proof is complete. ∎
Proof for Theorem 6.
$I_{p_{n}}(\theta^{*})$ $=\mathbb{E}_{k_{p_{n}}}\left[(\nabla_{\theta}\log k_{p_{n}}(\bm{x};\theta^{*}))(\nabla_{\theta}\log k_{p_{n}}(\bm{x};\theta^{*}))^{\top}\right]$
$=(1-p_{n})^{2}\int\frac{\left[\nabla_{\theta}f(\bm{x};\theta^{*})\right]\left[\nabla_{\theta}f(\bm{x};\theta^{*})\right]^{\top}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x.$
Let $H(\bm{x}):=\left[\nabla_{\theta}f(\bm{x};\theta^{*})\right]\left[\nabla_{\theta}f(\bm{x};\theta^{*})\right]^{\top}$ and let $[H(\bm{x})]_{i,j}$ denote its $(i,j)$-th entry. Consider the sets $A_{i,j}:=\{x:[H(\bm{x})]_{i,j}\geq 0\}$, we have
$0<\left|\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}\right|\leq\frac{[H(\bm{x})]_{i,j}}{(1-p_{n})f^{2}(\bm{x};\theta^{*})},\text{ for }x\in A_{i,j}$
and
$0<\left|\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}\right|\leq-\frac{[H(\bm{x})]_{i,j}}{(1-p_{n})f^{2}(\bm{x};\theta^{*})},\text{ for }x\in A_{i,j}^{c}.$
Note that $A_{i,j}$ and $A_{i,j}^{c}$ are both measurable with
$[I(\theta^{*})]_{i,j}=\int_{A_{i,j}\cup A^{c}_{i,j}}\frac{[H(\bm{x})]_{i,j}}{f^{2}(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x.$
Thus, apply dominated convergence theorem on both $A_{i,j}$ and $A_{i,j}^{c}$,
$\lim_{n\rightarrow\infty}\left[I_{p_{n}}(\theta^{*})\right]_{i,j}$ $\,=\lim_{n\rightarrow\infty}(1-p_{n})^{2}\left(\int_{A_{i,j}}\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x\right.$
$\,\qquad\qquad\qquad\qquad\left.+\int_{A_{i,j}^{c}}\frac{[H(\bm{x})]_{i,j}}{k_{p_{n}}(\bm{x};\theta^{*})f(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x\right)$
$\,=\int_{A_{i,j}\cup A_{i,j}^{c}}\frac{[H(\bm{x})]_{i,j}}{f^{2}(\bm{x};\theta^{*})}f(\bm{x};\theta^{*})\mathtt{d}x=[I(\theta^{*})]_{i,j}.$
Appendix B Proofs for Section 3
B.1 Results in Main Text
Definition 8 ($u$-cover under $\|\cdot\|_{\infty}$).
A $u$-cover of the set $\Omega$ is a set $\{\theta_{1},\dots,\theta_{n}\}\subset\mathbb{R}^{d}$ such that for any $\theta\in\Omega$, there exists $i\in\{1,\dots,n\}$ such that $\|\theta-\theta_{i}\|_{\infty}\leq u$.
Definition 9 (Covering Number).
The $u$-covering number of $\Omega$ under $\|\cdot\|_{\infty}$ is
$N(u,\Omega):=\min\{N\in\mathbb{N}:\exists\text{ a $u$-cover under $\|\cdot\|_{\infty}$ of size }N\}.$
Definition 10 ($u$-separated points under $\|\cdot\|_{\infty}$).
A set of points $\{\theta_{1},\dots,\theta_{k}\}\subset\Omega$ is $u$-separated if $\|\theta_{i}-\theta_{j}\|_{\infty}>u$, $\forall i\neq j$.
Definition 11 (Packing Number).
The $u$-packing number of $\Omega$ under $\|\cdot\|_{\infty}$ is
$D(u,\Omega):=\max\{N\in\mathbb{N}:\exists\text{ a set of $N$ points that is $u$-separate under $\|\cdot\|_{\infty}$}\}.$
B.1.1 Proof for Lemma 3
Proof.
The first half of this proof is based on several results commonly seen in Empirical processes literature (with some variation), see, for instance, Section 2.8 of Van der Vaart and Wellner (1996).
Recall that the (8) in Assumption 1 requires
$\int_{\phi_{n}^{2}/2^{10}}^{\phi_{n}}H^{1/2}_{[]}(u,\mathcal{F}(\Omega_{n}))\mathtt{d}u\leq c\sqrt{n}\phi_{n}^{2}.$
To deduce (8), we first prove the following bound on $N_{[]}$,
$N_{[]}(uM^{(n)}_{u},\mathcal{F}(\Omega_{n}))\leq C_{d}\text{Vol}(\tilde{\Omega}_{n})u^{-d}$ (25)
where $\tilde{\Omega}_{n}:=\{\theta\in\mathbb{R}^{d}:\exists\phi\in\Omega_{n},\|\theta-\phi\|_{\infty}\leq u/2\}$.
Let $\{\theta_{1},\dots,\theta_{k}\}$ be an $u$-cover of $\Omega_{n}$ with respect to $\|\cdot\|_{\infty}$. Consider the following functions, for $i=1,\dots,k$,
$\tilde{L}_{i}(\bm{x}):=\max\left\{\sqrt{f_{\theta_{i}}(\bm{x})}-uM^{(n)}_{\theta_{i},u}(\bm{x}),0\right\},\,\,\,\tilde{U}_{i}:=\sqrt{f_{\theta_{i}}(\bm{x})}+uM^{(n)}_{\theta_{i},u}(\bm{x}),$
and $L_{i}(\bm{x}):=\tilde{L}_{i}(\bm{x})^{2}$, $U_{i}:=\tilde{U}_{i}(\bm{x})^{2}$. Since
$|\sqrt{f_{\theta}(\bm{x})}-\sqrt{f_{\theta_{i}}(\bm{x})}|\leq M^{(n)}_{\theta_{i},u}(\bm{x})\|\theta-\theta_{i}\|_{\infty},$
we have, for any $\theta$ such that $\|\theta-\theta_{i}\|_{\infty}\leq u$,
$\tilde{L}_{i}(\bm{x})\leq\sqrt{f_{\theta}(\bm{x})}\leq\tilde{U}_{i}(\bm{x}).$
Moreover, recall that $M^{(n)}_{u}:=\sup_{\vartheta\in\Omega_{n}}\|M^{(n)}_{\vartheta,u}\|_{2}$,
$h^{2}(L_{i},U_{i})$ $=\frac{1}{2}\iint\left(\sqrt{L_{i}}-\sqrt{U_{i}}\right)^{2}\mathtt{d}\bm{x}$
$\leq\iint u^{2}M^{(n)}_{\theta_{i},u}(\bm{x})^{2}\mathtt{d}\bm{x}\leq(uM^{(n)}_{u})^{2}.$
Thus the set of $[L_{i},U_{i}]$ form a $uM^{(n)}_{u}$-bracketing for $\mathcal{F}(\Omega_{n})$ under the Hellinger distance $h(\cdot,\cdot)$, hence,
$N_{[]}(uM^{(n)}_{u},\mathcal{F}(\Omega_{n}))\leq N(u,\Omega_{n}).$
To complete (25), we recall that $N(u,\Omega_{n})\leq D(u,\Omega_{n})$ always holds. If $\{\theta_{1},\dots,\theta_{k}\}$, $k:=D(u,\Omega_{n})$, is a set of $u$-separated points, then the balls of radius $u/2$ centred at $\{\theta_{1},\dots,\theta_{k}\}$ are disjoint and contained in $\tilde{\Omega}_{n}:=\{\theta\in\Omega:\exists\theta^{\prime}\in\Omega_{n},\|\theta-\theta^{\prime}\|_{2}\leq u/2\}$. Thus $D(u,\Omega_{n})$ is upper bounded by the ratio between the volume of the space $\tilde{\Omega}_{n}$ and the $k$-collection ball of radius $u/2$. Let $v_{d}$ denote the volume of a unit ball and define $C_{d}:=2^{d}/v_{d}$ then
$N_{[]}(uM^{(n)}_{u},\mathcal{F}(\Omega_{n}))\leq N(u,\Omega)\leq D(u,\Omega)\leq C_{d}\text{Vol}(\tilde{\Omega}_{n})u^{-d}.$
Recall that $R_{n}$ is the radius of $\Omega_{n}$. $\phi_{n}$ is a decreasing sequence, let $\phi_{0}$ denote an upper bound for $\{\phi_{n}\}$ then $u\in(0,\phi_{0})$ and $\text{Vol}(\tilde{\Omega}_{n})=O((R_{n}+u/2)^{d})<O((R_{n}+\phi_{0})^{d})$, thus,
$H_{[]}(u,\mathcal{F}(\Omega_{n}))\leq\log C^{*}+d\log(R_{n}+\phi_{0})+d\log M_{u}^{(n)}-d\log u.$ (26)
Since $H_{[]}(u,\mathcal{F}(\Omega_{n})$ decreases with $u$ but $cnu^{2}$ increases with $u$, it suffices to check
$H_{[]}(\phi_{n}^{2}/2^{10},\mathcal{F}(\Omega_{n}))\leq c^{2}n\phi_{n}^{2}$
to prove that
$H_{[]}(u,\mathcal{F}(\Omega_{n}))\leq c^{2}n\phi_{n}^{2},\quad\forall u\in(\phi_{n}^{2}/2^{10},\phi_{n}).$
Due to the restriction that $n\phi_{n}^{2}\rightarrow\infty$, we have $\phi_{n}^{-1}=o(n^{1/2})$, thus every term in the right-hand side of (26) are of order $\log n$, apart from $\log M_{u}^{(n)}$. By assumption, there exists $q\in(0,1)$ such that $\forall u\in(0,\phi_{0})$,
$\frac{\log\log M_{u}^{(n)}}{\log n}<q,$
then $\log M_{u}^{(n)}=O(n^{q})$, and we can choose $\phi_{n}>n^{-\tfrac{1-q}{2}}$ such that
$H_{[]}(\phi_{n}^{2}/2^{10},\mathcal{F}(\Omega_{n}))\leq$ $\,\log C^{*}+d\log(R_{n}+\phi_{0})+d\log M_{\phi_{n}^{2}/2^{10}}^{(n)}-4d\log\phi_{n}$
$<$ $\,c^{2}n^{q}\leq c^{2}n\phi_{n}^{2}.$
for $n$ sufficiently large which implies (8).∎
B.1.2 Proof for Lemma 7
Proof.
Recall that
$\rho_{0,\alpha}(\theta^{*},\theta):=\frac{1}{\alpha}\int\left[\left(\frac{f(x;\theta^{*})}{f(x;\theta)}\right)^{\alpha}-1\right]f(x;\theta^{*})\mathtt{d}x.$
By the assumption that $f(x;\theta)$ has all directional derivatives in $\theta$ and the definition of $\tilde{\bm{\nabla}}$ from Definition 7,
$\left|\left(\frac{f(x;\theta^{*})}{f(x;\theta)}\right)^{\alpha}-1\right|\leq\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\alpha\left(\frac{f(x;\theta^{*})}{f(x;\theta)}\right)^{\alpha-1}\frac{f(x;\theta^{*})}{f(x;\theta)^{2}}\|\tilde{\bm{\nabla}}_{\theta}f(x;\tilde{\theta})\|_{2}\|\theta-\theta^{*}\|_{2}.$
Therefore, if $\|\theta-\theta^{*}\|_{2}\leq c_{r}$, then,
$\rho_{0,\alpha}(\theta^{*},\theta)\leq$ $\,\left(\max_{\theta,\tilde{\theta}\in\bar{S}(c_{r})}\int\left[\frac{f(x;\theta^{*})}{f(x;\theta)}\right]^{\alpha+1}\|\tilde{\bm{\nabla}}_{\theta}f(x;\tilde{\theta})\|_{2}\mathtt{d}x\right)\|\theta-\theta^{*}\|_{2}$
$\leq$ $\,M_{c_{r}}\|\theta-\theta^{*}\|_{2}.$
In other words, $\bar{S}(t_{n})\subseteq S_{0,\alpha}(M_{c_{r}}t_{n})\cap\bar{S}(c_{r})$, and $\bar{S}(t_{n}/M_{c_{r}})\subseteq S_{0,\alpha}(t_{n})$. ∎
B.1.3 Proof for Lemma 8
Proof.
Recall that $\bar{S}(r):=\{\theta\in\Omega:\|\theta^{*}-\theta\|_{2}\leq r\}$ and the prior distribution on $\theta$ is $\pi_{0}(\theta)\sim\mathcal{N}(0,I_{d})$. Let $\theta_{i}$ denote the $i$th dimension of vector $\theta$, then
$\mathbb{P}_{\pi_{0}}(\bar{S}(r))=\int_{\|\theta^{*}-\theta\|_{2}\leq r}(2\pi)^{-d/2}\exp\left(-\frac{1}{2}\sum_{i=1}^{d}\theta_{i}^{2}\right)\mathtt{d}\theta_{1}\cdots\mathtt{d}\theta_{d}.$
  1. 1.
    For the case when $r$ is small, let $w_{i}=\theta_{i}-\theta_{i}^{*}$, then $w_{i}\sim\mathcal{N}(\theta^{*}_{i},1)$. Thus $\sum_{i}w_{i}^{2}$ is a non-central $\chi^{2}$ distribution with $d$ degrees of freedom and non-centrality parameter $\lambda=\|\theta^{*}\|_{2}^{2}$ with probability density function given by (Patnaik, 1949)
    $f_{\chi^{2}}(x;d,\lambda)=\frac{1}{2}\exp\left(-\frac{x+\lambda}{2}\right)\left(\frac{x}{\lambda}\right)^{(d-2)/4}I_{d/2-1}(\sqrt{\lambda x})$
    where $I_{\nu}(x)$ is the modified Bessel function of the first kind,
    $I_{\nu}(x)=(x/2)^{\nu}\sum_{j=0}^{\infty}\frac{(x^{2}/4)^{j}}{j!\Gamma(\nu+j+1)}$
    Now,
    $\mathbb{P}_{\pi_{0}}(\bar{S}(r))=$ $\,\mathbb{P}\left(\chi^{2}(d,\lambda)\leq r^{2}\right)$
    $=$ $\,\int_{0}^{r^{2}}\frac{1}{2}\exp\left(-\frac{x+\lambda}{2}\right)\left(\frac{x}{\lambda}\right)^{(d-2)/4}I_{d/2-1}(\sqrt{\lambda x})\mathtt{d}x$
    $\geq$ $\,\int_{0}^{r^{2}}\frac{1}{2\Gamma(d/2)}\exp\left(-\frac{x+\lambda}{2}\right)\left(\frac{x}{\lambda}\right)^{(d-2)/4}\left(\frac{\sqrt{\lambda x}}{2}\right)^{d/2-1}\mathtt{d}x$
    $=$ $\,\frac{2^{-d/2}\exp(-\lambda/2)}{\Gamma(d/2)}\int_{0}^{r^{2}}x^{d/2-1}\exp(-x/2)\mathtt{d}x$
    $=$ $\,\frac{\exp(-\lambda/2)}{\Gamma(d/2)}\gamma(d/2,r^{2}/2).$
  2. 2.
    For the case where $r>\|\theta^{*}\|_{2}$, note that
    $\{\theta\in\Omega:\|\theta\|_{2}^{2}\leq(r-\|\theta^{*}\|_{2})^{2}\}\subseteq\{\theta\in\Omega:\|\theta^{*}-\theta\|_{2}^{2}\leq r^{2}\}.$
    Denote $R=r-\|\theta^{*}\|_{2}$,
    $\mathbb{P}_{\pi_{0}}(\bar{S}(r))\geq$ $\,\mathbb{P}\left(\chi^{2}(d,0)\leq R^{2}\right)$
    $=$ $\,\frac{1}{\Gamma(d/2)}\gamma(d/2,R^{2}/2).$
B.1.4 Proof for Proposition 4
Proof.
If Assumption 5 holds, then the same assumption holds for the set of density functions $k_{p}(\bm{x};\theta)$ for any $p\in(0,1)$. So it suffices to prove the proposition holds for $f$ given Assumption 5, then the same argument can be used to prove this proposition holds for any $k_{p}$, $p\in(0,1)$.
Recall that by Assumption 5, $\forall\psi>0$,
$\inf_{\|\theta-\theta^{*}\|_{2}>\psi}h(f(\cdot;\theta),f(\cdot;\theta^{*}))>0.$
Denote
$A(\psi):=\inf_{\|\theta-\theta^{*}\|_{2}>\psi}h(f(\cdot;\theta),f(\cdot;\theta^{*}))$
and $M=\sup_{\psi}A(\psi)>0$. Note that $A(\psi)$ is non-decreasing and right-continuous since the infimum is taken over the set excluding $\|\theta-\theta^{*}\|=\psi$. Then $\forall\phi<M$, $\exists\psi$ such that $\psi=\min\{\psi^{\prime}:\phi\leq A(\psi^{\prime})\}$. Therefore, $\forall\phi_{n}\rightarrow 0$, $\exists\psi_{n}$ non-increasing such that $\phi_{n}\leq A(\psi_{n})$ and we may choose $\psi_{n}:=\min\{\psi:\phi_{n}\leq A(\psi)\}$. Suppose for contradiction that $\psi_{n}\not\rightarrow 0$, then $\forall m>0$, $\exists N>0$ such that $\forall n>N$,
$\inf\{\psi:\phi_{n}\leq A(\psi)\}>m.$
Thus, $\forall\psi^{\prime}\leq m$, we have $A(\psi^{\prime})<\phi_{n}$, but since $\phi_{n}\rightarrow 0$, we have
$\inf_{\|\theta-\theta^{*}\|_{2}>m}h(f(\cdot;\theta),f(\cdot;\theta^{*}))=0,$
which contradicts the assumption. Hence $\psi_{n}\rightarrow 0$ and the proof is completed.
B.2 Proofs Related to Lemma 9
Recall that $d(x;\theta):=k_{p}(x;\theta)/k_{p}(x;\theta^{*})$ where $\theta^{*}$ is the true paramter. For $\tau>0$ to be chosen later, let
$\tilde{Z}_{\theta}(\bm{x}):=\max\{-\tau,\log d(\bm{x},\theta)\}$
and
$\nu_{n}(f):=n^{-1/2}\sum_{i=1}^{n}\bm{X}_{i}-\mathbb{E}\bm{X}_{i}.$
Theorem 16 (One-sided Bernstein’s Inequality).
Let $X_{1},\dots,X_{n}$ be i.i.d. random variables, then the following holds:
  • If $|X_{i}-\mathbb{E}X_{i}|\leq T,\forall i\in\{1,\dots,n\}$ and $\text{Var}(X_{i})\leq v$, then
    $\mathbb{P}^{*}\left(n^{-1/2}\sum_{i=1}^{n}(X_{i}-\mathbb{E}X_{i})>M\right)\leq\exp\left(-\frac{M^{2}}{2(v+MT/3n^{1/2})}\right);$ (27)
  • If $\mathbb{E}|X_{i}|^{j}\leq j!b^{j-2}v/2$ for any $j\geq 2$, then
    $\mathbb{P}^{*}\left(n^{-1/2}\sum_{i=1}^{n}(X_{i}-\mathbb{E}X_{i})>M\right)\leq\exp\left(-\frac{M^{2}}{2(v+bM/n^{1/2})}\right)$ (28)
where $\mathbb{P}^{*}$ denotes the outer measure with respect to the measure of $\bm{X}_{i}$s.
Proof.
See Bennett (1962).∎
Lemma 17 (Lemma 7 (Wong and Shen, 1995)).
Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be i.i.d. random variables following density $k_{p}(\cdot;\theta^{*})$. For any $t>0$, $10<k<$ 1, $M>0$, let
$\Psi(M,t^{2},n):=\frac{M^{2}}{4(8c_{0}t^{2}+M/n^{1/2})}$
where
$c_{0}:=\frac{\exp(\tau/2)-1-\tau/2}{(1-\exp(-\tau/2))^{2}},\quad M:=kn^{1/2}t^{2},$
and $\tau>0$ is a constant to be chosen. Let
$s(t):=\frac{kM}{16\sqrt{2n}\exp(\tau/2)}=\frac{k^{2}t^{2}}{2^{4.5}\exp(\tau/2)}.$
If
$\int_{s(t)/4}^{t}H_{[]}^{1/2}(u,\mathcal{K}_{p}(\Omega_{n}))\mathtt{d}u\leq\frac{\sqrt{3}k^{3/2}M}{2^{10}\exp(\tau/2)}=\frac{\sqrt{3}k^{5/2}n^{1/2}t^{2}}{2^{10}\exp(\tau/2)}$ (29)
then
$\mathbb{P}^{*}\left(\sup_{\begin{subarray}{c}h(\theta,\theta^{*})\leq t\\ \theta\in\Omega_{n}\end{subarray}}\nu(\tilde{Z}_{\theta})\geq M\right)\leq(1+N(t))\exp(-(1-k)\Psi(M,t^{2},n))$
where
$N(t):=\min\{x\in\mathbb{N}:2^{x}s(t)>t\}$
Proof.
The proof follows from a similar chaining argument as in Shen and Wong (1994). In this proof only, we will also consider the bracketing entropy computed in $L_{2}$-norm denoted by $H_{[]}(u,\mathcal{F},\|\cdot\|_{2})$. To make further distinction, we will denote the Hellinger bracketing entropy by $H_{[]}(u,\mathcal{F},h)$.
Recall that for $\theta\in\Omega_{n}$, we define $\tilde{Z}_{\theta}(\bm{x}):=\max\{-\tau,\log d(\bm{x},\theta)\}$. Let $\mathcal{G}:=\mathcal{K}_{p}(\Omega_{n})$ denote the space of density functions of interest, and let $\tilde{\mathcal{Z}}:=\{\tilde{Z}_{\theta}:\theta\in\Omega_{n}\}$ be the space of lower-truncated log-likelihood ratios. Note that the two spaces are linked, in the sense that the Hellinger bracketing entropy of $\mathcal{G}$ is related to the $L_{2}$-bracketing entropy of $\tilde{\mathcal{Z}}$ by
$H_{[]}(2\sqrt{2}\exp(\tau/2)\epsilon,\tilde{\mathcal{Z}},\|\cdot\|_{2})\leq H_{[]}(\epsilon,\mathcal{G},h),$
see Lemma 3 of Wong and Shen (1995) or noting that
$\|\tilde{Z}_{\theta_{1}}-\tilde{Z}_{\theta_{2}}\|_{2}^{2}\leq 4\exp(\tau)\|k_{p}^{1/2}(\cdot;\theta_{1})-k_{p}^{1/2}(\cdot;\theta_{2})\|_{2}^{2}=8\exp(\tau)h^{2}(\theta_{1},\theta_{2}).$ (30)
Provided that (which we will prove directly later)
$H_{[]}(t,\mathcal{G},h)\leq\frac{k}{4}\Psi(M,t^{2},n),$ (31)
we can construct a sequence of bracketings on the space $\mathcal{G}$ and $\tilde{\mathcal{Z}}$. Let $t\geq\delta_{0}>\delta_{1}>\cdots>\delta_{N}>0$, and $\mathcal{F}_{0},\dots\mathcal{F}_{N}$ denote the set of Hellinger bracketings of $\mathcal{G}$ with bracket size $\delta_{0},\dots,\delta_{N}$ respectively. We may choose $\delta_{j}$ such that
$\delta_{0}:=\max\{\delta\leq t:H_{[]}(\delta,\mathcal{G},h)\leq\tfrac{k}{4}\Psi(M,t^{2},n)\}$
and for $j=0,\dots,N-1$,
$\delta_{j+1}:=\max\{s(t),\sup\{\delta\leq\delta_{j}/2:H_{[]}(\delta,\mathcal{G},h)\geq 4H_{[]}(\delta_{j},\mathcal{G},h)\}\},$
with $N:=\min\{j:\delta_{j}\leq s\}$. Furthermore, let
$\tilde{\delta}_{0}:=\left(\sup\{\tilde{\delta}\leq 2\sqrt{2}\delta_{0}\exp(\tau/2):H_{[]}(\tilde{\delta},\tilde{\mathcal{Z}},\|\cdot\|_{2})=H_{[]}(\delta_{0},\mathcal{G},h)\}\right)^{-}$
and $\tilde{\delta}_{j}:=2\sqrt{2}\delta_{j}\exp(\tau/2)$, then
$H_{[]}(\tilde{\delta}_{j},\tilde{\mathcal{Z}},\|\cdot\|_{2})\leq H_{[]}(\delta_{j},\mathcal{G},h)$
Note that the entropy function $H_{[]}$ is decreasing and right-continuous. We want to evaluate $H_{[]}(\tilde{\delta}_{0},\tilde{\mathcal{Z}},\|\cdot\|_{2})$ at the left limit $(\tilde{\delta}_{0})^{-}$ if a jump happens at $\tilde{\delta}_{0}$, so that $H_{[]}(\tilde{\delta}_{0},\tilde{\mathcal{Z}},\|\cdot\|_{2})=H_{[]}(\delta_{0},\mathcal{G},h)$.
The space $\tilde{\mathcal{Z}}$ can be obtained by transforming elements of $\mathcal{G}$ through mapping $f\mapsto\max\{-\tau,\log(f/k_{p}(\cdot;\theta^{*}))\}$, we may assume the choice of bracketing in one space induces another for the case $j=0$.
For each $\theta\in\Omega_{n}$, there exists a bracket $[\tilde{Z}^{L}_{j,\theta},\tilde{Z}^{U}_{j,\theta}]$ for every $j$ such that $\tilde{Z}^{L}_{j,\theta}\leq\tilde{Z}_{\theta}\leq\tilde{Z}^{U}_{j,\theta}$ and $\|\tilde{Z}^{L}_{j,\theta}-\tilde{Z}^{U}_{j,\theta}\|_{2}\leq\tilde{\delta}_{j}$. Define a sequence of shrinking envelop functions for each $\theta$ by
$u_{j}(\theta):=\min_{i\leq j}\tilde{Z}^{U}_{i,\theta},\quad l_{j}(\theta):=\max_{i\leq j}\tilde{Z}^{L}_{i,\theta}.$
Then $l_{j}(\theta)\leq\tilde{Z}_{\theta}\leq u_{j}(\theta)$, and $\|u_{j}-l_{j}\|_{2}\leq\tilde{\delta}_{j}$. For notional simplicity, we hide the dependency of $u_{j}(\theta),l_{j}(\theta)$ on $\theta$ and write them directly as $u_{j}$ and $l_{j}$ as functions of $\bm{X}_{i}$. Finally, define sequences $a_{1}>\cdots>a_{N}$, $\eta_{0},\dots,\eta_{N}$ by
$a_{j}:=\frac{\sqrt{n}\tilde{\delta}_{j-1}^{2}}{\eta_{j-1}},\quad\eta_{j}:=\frac{4\tilde{\delta}_{j}}{\sqrt{3k}}\left(\sum_{i\leq j+1}H_{[]}(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})\right)^{1/2},\quad\eta_{N}:=0.$ (32)
We may decompose $\tilde{Z}_{\theta}$ on a partition on the sample space $B_{0},\dots,B_{N}$
$\tilde{Z}_{\theta}=$ $\,u_{0}+\sum_{j=0}^{N}(u_{j}I_{B_{j}}-u_{0}I_{B_{j}})+\left(\tilde{Z}_{\theta}-\sum_{j=0}^{N}u_{j}I_{B_{j}}\right)$
$=$ $\,u_{0}+\sum_{j=1}^{N}\sum_{i=1}^{j}(u_{i}-u_{i-1})I_{B_{j}}+\sum_{j=0}^{N}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}$
$=$ $\,u_{0}+\sum_{j=1}^{N}(u_{j}-u_{j-1})I_{\cup_{i\geq j}B_{i}}+\sum_{j=0}^{N}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}$ (33)
where
$B_{0}:=\{(u_{0}-l_{0})\geq a_{1}\},$
$B_{j}:=\{(u_{j}-l_{j})\geq a_{j+1},\,\,\,(u_{i}-l_{i})<a_{i+1},i=1,\dots,j-1\},$
$B_{N}:=\left(\cup_{j=0}^{N-1}B_{j}\right)^{c}.$
Note that
$\sum_{j=0}^{N}\eta_{j}=$ $\,\sum_{j=0}^{N-1}4\tilde{\delta}_{j}(3k)^{-1/2}\left(\sum_{i\leq j+1}H_{[]}(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})\right)^{1/2}$
$\leq$ $\,\frac{4\sqrt{2}}{\sqrt{3k}}\sum_{j=0}^{N-1}\tilde{\delta}_{j}(H(\delta_{j+1},\mathcal{G},h))^{1/2}$
$\leq$ $\,\frac{2^{7}}{\sqrt{3k}}\exp(\tau/2)\int_{s/4}^{t}H_{[]}^{1/2}(u,\mathcal{G},h)\mathtt{d}u$
$\leq$ $\,\frac{kM}{8}$
where the second-last step is by Lemma 3.1 of Alexander (1984) and the last step is by (29). Thus, through the decomposition of (33), we note
$\mathbb{P}^{*}\left(\sup_{\begin{subarray}{c}h(\theta,\theta^{*})\leq t\\ \theta\in\Omega_{n}\end{subarray}}\nu(\tilde{Z}_{\theta})\geq M\right)\leq$ $\,\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(u_{0})>(1-\tfrac{3k}{8})M\right)$
$+$ $\,\sum_{j=0}^{N-1}\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(u_{j+1}-u_{j})I_{i\geq j+1}B_{i}>\eta_{j}\right)$
$+$ $\,\sum_{j=0}^{N-1}\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}>\eta_{j}\right)$
$+$ $\,\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{N})I_{B_{N}}>\tfrac{kM}{8}+\eta_{N}\right)$
$\leq$ $\,\mathbb{P}_{1}+\mathbb{P}_{2}+\mathbb{P}_{3}+\mathbb{P}_{4}$
where
$\mathbb{P}_{1}:=$ $\,|\mathcal{F}_{0}|\sup_{h(\theta,\theta^{*})\leq t}\mathbb{P}(\nu_{n}(u_{0})>(1-\tfrac{3k}{8})M)$
$\mathbb{P}_{2}:=$ $\,\sum_{j=0}^{N-1}\prod_{l_{1}=0}^{j}|\mathcal{F}_{l_{1}}|\prod_{l_{2}=0}^{j+1}|\mathcal{F}_{l_{2}}|\sup_{h(\theta,\theta^{*})\leq t}\mathbb{P}(\nu_{n}(u_{j+1}-u_{j})I_{\cup_{i\geq j+1}B_{i}}>\eta_{j})$
$\mathbb{P}_{3}:=$ $\,\sum_{j=0}^{N-1}\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}>\eta_{j}\right)$
$\mathbb{P}_{4}:=$ $\,\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{N})I_{B_{N}}>\tfrac{kM}{8}+\eta_{N}\right).$
Before bounding $\mathbb{P}_{i}$, we can quickly check (31). Note that Hellinger distance has an upper bound of 1, so we may assume $t\leq 1$. Then by definition of $s(t)$, $s(t)\leq t/2\sqrt{2}\leq t/2$, so by $H_{[]}(u,\mathcal{G},h)$ decreasing,
$H_{[]}(t,\mathcal{G},h)\leq\left(\frac{1}{t-t/8}\int_{s/4}^{t}H_{[]}^{1/2}(u,\mathcal{G},h)\mathtt{d}u\right)^{2}\leq\frac{3k^{5}nt^{2}}{2^{14}\cdot 49\exp(\tau)}\leq\frac{k^{3}nt^{2}}{2^{7}c_{0}+16k}=\frac{k}{4}\Psi(M,t^{2},n).$
for any $32^{7}c_{0}+16k\leq 2^{14}\cdot 49/$ 3, which can be easily satisfied by a wide range of choices, including the value choice we will use in the proof for main results.
  • To bound $\mathbb{P}_{1}$, we use (28). By Lemma 5 (Wong and Shen, 1995), we have
    $\mathbb{E}|u_{0}|^{j}\leq j!2^{j}c_{0}\|w^{1/2}(\cdot)-k_{p}^{1/2}(\cdot,\theta^{*})\|_{2}^{2},\,\,\forall j\geq 2,$
    where $w(x):=\exp(u_{0}(x))k_{p}(x,\theta^{*})$. Note that
    $\|w^{1/2}(\cdot)-k_{p}^{1/2}(\cdot,\theta^{*})\|_{2}^{2}\leq\|w^{1/2}-k_{p}^{1/2}(,\theta)\|_{2}^{2}+h^{2}(\theta,\theta^{*})\leq 2t^{2}.$
    Thus, recall Bernstein’s inequality (28) and letting $b:=2$, $v:=8c_{0}\|w^{1/2}(\cdot)-k_{p}^{1/2}(\cdot,\theta^{*})\|_{2}^{2}$,
    $\mathbb{P}_{1}\leq$ $\,|\mathcal{F}_{0}|\mathbb{P}(\nu_{n}(u_{0})>(1-\tfrac{3k}{8})M)$
    $\leq$ $\,\exp\left(H_{[]}(\delta_{0},\mathcal{G},h)-\Psi((1-\tfrac{3k}{8})M,t^{2},n)\right)$
    $\leq$ $\,\exp\left(\tfrac{k}{4}\Psi(M,t^{2},n)-(1-\tfrac{3k}{8})^{2}\Psi(M,t^{2},n)\right)$
    $\leq$ $\,\exp\left(-(1-k)\Psi(M,t^{2},n)\right),$
    where we use (31) to bound $H_{[]}(\delta_{0},\mathcal{G},h)$ at the second step.
  • $\mathbb{P}_{2}:=\sum_{j=0}^{N-1}\prod_{l_{1}=0}^{j}|\mathcal{F}_{l_{1}}|\prod_{l_{2}=0}^{j+1}|\mathcal{F}_{l_{2}}|\sup_{h(\theta,\theta^{*})\leq t}\mathbb{P}(\nu_{n}(u_{j+1}-u_{j})I_{\cup_{i\geq j+1}B_{i}}>\eta_{j})$
    By the definition of $B_{j}$, on the set $\cup_{i\geq j+1}B_{j}$, we have $-a_{j+1}<l_{j}-u_{j}<u_{j+1}-u_{j}<0$, thus
    $\text{Var}((u_{j+1}-u_{j})I_{\cup_{i\geq j+1}B_{i}})\leq$ $\,\mathbb{E}[(u_{j+1}-u_{j})^{2}I_{\cup_{i\geq j+1}B_{i}}]$
    $\leq$ $\,\mathbb{E}[(u_{j+1}-u_{j})^{2}]\leq\tilde{\delta}_{j}^{2}.$
    Thus, by the one-sided Bernstein’s inequality for bounded random variables (27),
    $\mathbb{P}(\nu_{n}(u_{j+1}-u_{j})I_{\cup_{i\geq j+1}B_{i}}>\eta_{j})\leq\exp\left(-\frac{\eta_{j}^{2}}{2(\tilde{\delta}_{j}^{2}+a_{j+1}\eta_{j}/3n^{1/2})}\right).$
    For $j=0$, note that
    $(u_{1}-u_{0})I_{\cup_{j\geq 1}B_{j}}-\mathbb{E}[(u_{1}-u_{0})I_{\cup_{j\geq 1}B_{j}}]\leq\left|\mathbb{E}[(u_{1}-u_{0})I_{\cup_{j\geq 1}B_{j}}]\right|\leq\tilde{\delta}_{0},$
    hence by (27) again,
    $\mathbb{P}(\nu_{n}(u_{1}-u_{0})I_{\cup_{i\geq 1}B_{i}}>\eta_{0})\leq\exp\left(-\frac{\eta_{0}^{2}}{2(\tilde{\delta}_{0}^{2}+\tilde{\delta}_{0}\eta_{0}/3n^{1/2})}\right).$
    Recall from (32), for $j=0,\dots,N-1$,
    $\eta_{j}:=4\tilde{\delta}_{j}k^{-1/2}\left(\sum_{i\leq j+1}H_{[]}(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})\right)^{1/2}$
    Now, by $\eta_{0}\leq\sum\eta_{j}\leq kM/8=k^{2}n^{1/2}t^{2}/8$, we have
    $\frac{\eta_{0}^{2}}{2(\tilde{\delta}_{0}^{2}+\tilde{\delta}_{0}\eta_{0}/3n^{1/2})}\geq$ $\,\frac{3\eta_{0}^{2}}{2(3\tilde{\delta}_{0}^{2}+\tilde{\delta}_{0}k^{2}t^{2}/8)}$
    $\geq$ $\,\frac{16}{6+k^{2}t^{2}/4\tilde{\delta}_{0}}\cdot\frac{\sum_{j\leq 1}H(\tilde{\delta}_{j},\tilde{\mathcal{Z}},\|\cdot\|_{2})}{k}$
    $\geq$  $\,2\sum_{j\leq 1}\frac{H(\tilde{\delta}_{j},\tilde{\mathcal{Z}},\|\cdot\|_{2})}{k},$
    where the last step is due to $\tilde{\delta}_{0}\geq 2\sqrt{2}\exp(\tau/2)s=k^{2}t^{2}/8$. For $j\geq 1$,
    $\frac{\eta_{j}^{2}}{2(\tilde{\delta}_{j}^{2}+a_{j+1}\eta_{j}/3n^{1/2})}\geq$ $\,\frac{3\eta_{j}^{2}}{6\tilde{\delta}_{j}^{2}+2a_{j+1}\eta_{j}/\sqrt{n}}$
    $\geq$ $\,\frac{3\eta_{j}^{2}}{8\tilde{\delta}_{j}^{2}}\geq 2\sum_{i\leq j+1}\frac{H(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})}{k}$
    which follows directly from the definition of $a_{j+1}$ and $\eta_{j}$. Now,
    $\mathbb{P}_{2}:=$ $\,\sum_{j=0}^{N-1}\prod_{l_{1}=0}^{j}|\mathcal{F}_{l_{1}}|\prod_{l_{2}=0}^{j+1}|\mathcal{F}_{l_{2}}|\sup_{h(\theta,\theta^{*})\leq t}\mathbb{P}(\nu_{n}(u_{j+1}-u_{j})I_{\cup_{i\geq j+1}B_{i}}>\eta_{j})$
    $\leq$ $\,\exp\left(2\sum_{j\leq 1}H(\tilde{\delta}_{j},\tilde{\mathcal{Z}},\|\cdot\|_{2})-\frac{\eta_{0}^{2}}{2(\tilde{\delta}_{0}^{2}+\tilde{\delta}_{0}\eta_{0}/3n^{1/2})}\right)$
    $+$ $\,\sum_{j=1}^{N-1}\exp\left(2\sum_{i\leq j+1}H(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})-\frac{\eta_{j}^{2}}{2(\tilde{\delta}_{j}^{2}+a_{j+1}\eta_{j}/3n^{1/2})}\right)$
    $\leq$ $\,\sum_{j=0}^{N-1}\exp\left(-2\frac{1-k}{k}\sum_{i\leq j+1}H(\tilde{\delta}_{i},\tilde{\mathcal{Z}},\|\cdot\|_{2})\right)$
    $\leq$ $\,\sum_{j=0}^{N-1}\exp(-2(1-k)4^{j}\Psi(M,t^{2},n))\leq N(t)\exp(-(1-k)\Psi(M,t^{2},n)).$ (34)
  • For $\mathbb{P}_{3}$ and $\mathbb{P}_{4}$, we try to show that they are both 00.
    $\mathbb{P}_{3}:=\sum_{j=0}^{N-1}\mathbb{P}^{*}\left(\sup_{h(\theta,\theta^{*})\leq t}\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}>\eta_{j}\right)$
    Note that by the definition of $\nu_{n}$
    $\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}\leq n^{-1/2}\sum_{j=1}^{n}(\tilde{Z}_{\theta}(\bm{X}_{i})-u_{j}(\bm{X}_{i}))I_{B_{j}}+n^{1/2}\sup_{h(\theta,\theta^{*})\leq t}\mathbb{E}[(u_{j}-l_{j})I_{B_{j}}].$
    On $B_{j}$, $u_{j}-l_{j}\geq a_{j+1}$, thus $\mathbb{P}(B_{j})\leq\mathbb{E}[(u_{j}-l_{j})^{2}]/a_{j+1}^{2}\leq\tilde{\delta}_{j}^{2}/a_{j+1}^{2}$, and
    $\sup_{h(\theta,\theta^{*})\leq t}\mathbb{E}[(u_{j}-l_{j})I_{B_{j}}]=$ $\,\sup_{h(\theta,\theta^{*})\leq t}\left(\mathbb{E}[(u_{j}-l_{j})^{2}]\mathbb{E}[I_{B_{j}}]\right)^{1/2}$
    $\leq$ $\,\frac{\tilde{\delta}_{j}^{2}}{a_{j+1}}\leq\frac{\eta_{j}}{\sqrt{n}}.$
    Hence $\nu_{n}(\tilde{Z}_{\theta}-u_{j})I_{B_{j}}\leq\eta_{j}$, and $\mathbb{P}_{3}=0$.
  • Similarly,
    $\nu_{n}(\tilde{Z}_{\theta}-u_{N})I_{B_{N}}\leq$ $\,n^{1/2}\sup_{h(\theta,\theta^{*})\leq t}\mathbb{E}[(u_{N}-l_{N})I_{B_{N}}]$
    $\leq$ $\,n^{1/2}\tilde{\delta}_{N}=2\sqrt{2}n^{1/2}\exp(\tau/2)s(t)=\frac{kM}{8},$
    therefore, $\mathbb{P}_{4}=0$.
Sum through $\mathbb{P}_{1}$ to $\mathbb{P}_{4}$ completes the proof. ∎
Corollary 18.
With the same definitions as in Lemma 17, if in addition that
$\exp(-(1-k)\Psi(M,t^{2},n))\leq\frac{1}{2},$ (35)
then
$\mathbb{P}^{*}\left(\sup_{\begin{subarray}{c}h(\theta,\theta^{*})\leq t\\ \theta\in\Omega_{n}\end{subarray}}\nu(\tilde{Z}_{\theta})\geq M\right)\leq 2\exp(-(1-k)\Psi(M,t^{2},n)).$
Proof.
With the same proof as in Lemma 17, but (34) can be instead bounded by
$\,\sum_{j=0}^{N-1}\exp(-2(1-k)4^{j}\Psi(M,t^{2},n))$
$\leq$ $\,\exp(-(1-k)\Psi(M,t^{2},n))^{2}\left(\sum_{j=0}^{\infty}\exp(-(1-k)4j\Psi(M,t^{2},n)\right)$
$\leq$ $\,\exp(-(1-k)\Psi(M,t^{2},n))\times\frac{1}{2}\times\frac{16}{15}$
$\leq$ $\,\exp(-(1-k)\Psi(M,t^{2},n)).$
due to the additional assumption (35). ∎
Proof for Lemma 9.
By the choice of $\tau$ and $k$ we use later, we see that $t^{2}/2^{10}<s(t)$, thus (20) implies (29), treating $\phi_{n}$ as $t$.If (29) holds for some $t>0$, then it holds for any $\tilde{t}\geq t$. Thus if (20) holds, we may apply the result of Lemma 17 on any $\tilde{t}\geq t$, then
$\mathbb{P}^{*}\left(\sup_{\begin{subarray}{c}h(\theta,\theta^{*})\leq\tilde{t}\\ \theta\in\Omega_{n}\end{subarray}}\nu(\tilde{Z}_{\theta})\geq M(\tilde{t})\right)\leq$ $\,(1+N(\tilde{t}))\exp(-(1-k)\Psi(M,\tilde{t}^{2},n))$
$\leq$ $\,(1+N(t))\exp\left(-\frac{(1-k)k^{2}}{2^{5}c_{0}+4k}n\tilde{t}^{2}\right)$
From Lemma 4 (Wong and Shen, 1995),
$\mathbb{E}\tilde{Z}_{\theta}\leq-2(1-\varrho)h^{2}(\theta,\theta^{*})$
where $\varrho:=2\exp(-\tau/2)/(1-\exp(-\tau/2))^{2}$. Let $A_{\tilde{t}}:=\{\theta\in\Omega_{n}:\tilde{t}\leq h(\theta,\theta^{*})\leq\sqrt{2}\tilde{t}\}$, we have for any $\tilde{t}>t$,
$\left\{\sup_{A_{\tilde{t}}}\sum_{i=1}^{n}\log d(\bm{x}_{i},\theta)\geq-2(1-\varrho-k/2)n\tilde{t}^{2}\right\}\subseteq\left\{\sup_{A_{\tilde{t}}}\nu_{n}(\tilde{Z}_{\theta})\geq k\sqrt{n}\tilde{t}^{2}\right\}.$
Therefore
$\,\mathbb{P}\left(\sup_{A_{\tilde{t}}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq\exp\left(-2(1-\varrho-k/2)n\tilde{t}^{2}\right)\right)$ (36)
$\leq$ $\,\mathbb{P}\left(\sup_{A_{\tilde{t}}}\nu_{n}(\tilde{Z}_{\theta})\geq k\sqrt{n}\tilde{t}^{2}\right)$
$\leq$ $\,(1+N(t))\exp\left(-\frac{(1-k)k^{2}}{2^{5}c_{0}+4k}n\tilde{t}^{2}\right).$
Let $L(t):=\min\{x\in\mathbb{N}:2^{x+1}t^{2}>1\}$, then
$\,\mathbb{P}^{*}\left(\sup_{A^{c}\cap\Omega_{n}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq\exp\left(-2(1-\varrho-k/2)nt^{2}\right)\right)$
$=$ $\,\sum_{j=0}^{L}\mathbb{P}^{*}\left(\sup_{2^{j}t^{2}\leq h(\theta,\theta^{*})<2^{j+1}t^{2}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq\exp\left(-2(1-\varrho-k/2)nt^{2}\right)\right)$
$\leq$ $\,\sum_{j=0}^{L}\mathbb{P}^{*}\left(\sup_{2^{j}t^{2}\leq h(\theta,\theta^{*})<2^{j+1}t^{2}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq\exp\left(-2(1-\varrho-k/2)n2^{j+1}t^{2}\right)\right)$
$\leq$ $\,\sum_{j=0}^{L}(1+N(t))\exp\left(-\frac{(1-k)k^{2}}{2^{5}c_{0}+4k}n2^{j}t^{2}\right)$ (37)
$\leq$ $\,(L(t)+1)(N(t)+1)\exp\left(-\frac{(1-k)k^{2}}{2^{5}c_{0}+4k}nt^{2}\right).$
To derive the choice of constants, we follow the suggestion of Wong and Shen (1995) and choose $k=2/3$, $\exp(\tau/2)=5$, then $c_{0}=25(4-\tau/2)/16\approx 3.74$, $c_{1}=1/12$, $c_{2}=4/27(32c_{0}+8/3)\approx 1.212\times 10^{-3}$ and $c=2^{-7.5}*3^{-1.5}/5\approx 2^{-13}$. Also, for uniform bound, $N\leq\left\lceil 8+\log_{2}(t^{-1})\right\rceil$ and $L=\left\lceil 2\log_{2}(t^{-1})-1\right\rceil$, in which case $(1+L)(1+N)\leq 20\max\{-\log_{2}(t),(\log_{2}(t))^{2}\}$.
For the last result, note that $\exp(-c_{2}nt^{2})\leq\frac{1}{2}$ implies (31), thus by Corollary 18, we can replace $(1+N)$ by $2$ in (37) and get
$\sum_{j=0}^{L}2\exp\left(-c_{2}n2^{j}t^{2}\right)\leq 4\exp\left(-c_{2}nt^{2}\right).$
Finally, relabel $t$ as $\phi_{n}$ to get the expression in the statement of the Lemma. ∎
B.3 Remaining Proofs for Section 3
B.3.1 Proof for Lemma 10
Proof for Lemma 10.
It is enough to check that a $u$-bracket for $\mathcal{K}_{p^{\prime}}(A)$ is also a $u\sqrt{\frac{1-p}{1-p^{\prime}}}$-bracketing for $\mathcal{K}_{p}(A)$, where $1>p\geq p^{\prime}$. Suppose that $\{[L^{\prime}_{i},U^{\prime}_{i}],i=1,\dots,N\}$ is a $u$-bracketing for $\mathcal{K}_{p^{\prime}}(A)$. Pick an arbitrary $\theta\in A$, there exists $i\in\{1,\dots,N\}$ such that the $u$-bracket $[L^{\prime}_{i},U^{\prime}_{i}]$ forms an envelope for $k_{p^{\prime}}(\bm{x};\theta)$, i.e.,
$L^{\prime}_{i}(\bm{x})\leq(1-p^{\prime})f(\bm{x};\theta)+p^{\prime}g(\bm{x})\leq U^{\prime}_{i}(\bm{x}),$
for which $h(L^{\prime}_{i},U^{\prime}_{i})\leq u$. Note that by direct multiplication and addition to the preceding inequality, we can find $L_{i}$ and $U_{i}$ such that
$L_{i}(\bm{x}):=\frac{1-p}{1-p^{\prime}}L^{\prime}_{i}(\bm{x})+\frac{p-p^{\prime}}{1-p^{\prime}}g(\bm{x})\leq k_{p}(\bm{x};\theta)\leq\frac{1-p}{1-p^{\prime}}U^{\prime}_{i}(\bm{x})+\frac{p-p^{\prime}}{1-p^{\prime}}g(\bm{x})=:U_{i}(\bm{x}).$
Thus
$h^{2}(L_{i},U_{i})=$ $\,\frac{1}{2}\int\left(\sqrt{L_{i}(\bm{x})}-\sqrt{U_{i}(\bm{x})}\right)^{2}\mathtt{d}\bm{x}$
$\leq$ $\,\frac{1}{2}\int\left(\sqrt{\frac{1-p}{1-p^{\prime}}L^{\prime}_{i}(\bm{x})}-\sqrt{\frac{1-p}{1-p^{\prime}}U^{\prime}_{i}(\bm{x})}\right)^{2}\mathtt{d}\bm{x}$
$=$ $\,\frac{1-p}{1-p^{\prime}}\cdot\frac{1}{2}\int\left(\sqrt{L^{\prime}_{i}(\bm{x})}-\sqrt{U^{\prime}_{i}(\bm{x})}\right)^{2}\mathtt{d}\bm{x}$
$=$ $\,\frac{1-p}{1-p^{\prime}}u^{2}$
where the second inequality is by noting that for any $a\geq b\geq 0,c\geq 0$, $\sqrt{a+c}-\sqrt{b+c}\leq\sqrt{a}-\sqrt{b}$. Thus $\{[L_{i},U_{i}],i=1,\dots,N\}$ is an $u\sqrt{\frac{1-p}{1-p^{\prime}}}$-bracketing for $\mathcal{K}_{p}(A)$.
For the second statement, note that $H_{[]}(u,\mathcal{F})$ is decreasing as $u$ increases and $\mathcal{F}(A)=\mathcal{K}_{0}(A)$. ∎
B.3.2 Proof for Lemma 12
Proof.
Recall that
$\rho_{p,\alpha}(\theta^{*},\theta):=\frac{1}{\alpha}\int\left[\left(\frac{k_{p}(\bm{x};\theta^{*})}{k_{p}(\bm{x};\theta)}\right)^{\alpha}-1\right]k_{p}(\bm{x};\theta)\mathtt{d}x.$
Note that the above expression is a special case of the $f$-divergence with convex function
$f(u)=\frac{1}{\alpha}\left(u^{-\alpha}-1\right),$
and to emphasize this fact, we will rewrite the expression as
$\rho_{\alpha}\left(k_{p}(\bm{x};\theta^{*})\|k_{p}(\bm{x};\theta)\right):=\frac{1}{\alpha}\int\left[\left(\frac{k_{p}(\bm{x};\theta^{*})}{k_{p}(\bm{x};\theta)}\right)^{\alpha}-1\right]k_{p}(\bm{x};\theta)\mathtt{d}x.$
By joint convexity of $f$-divergence,
$\rho_{p,\alpha}(\theta^{*},\theta)$ $\,=\rho_{\alpha}\left.\Bigl{(}(1-p)f(\bm{x};\theta^{*})+p\,g(\bm{x})\right\|(1-p)f(\bm{x};\theta)+p\,g(\bm{x})\Bigr{)}$
$\,\leq(1-p)\rho_{\alpha}\left.\bigl{(}f(\bm{x};\theta^{*})\right\|f(\bm{x};\theta)\bigr{)}+p\,\rho_{\alpha}\left.\bigl{(}g(\bm{x})\right\|g(\bm{x})\bigr{)}$
$\,=(1-p)\rho_{\alpha}\left.\bigl{(}f(\bm{x};\theta^{*})\right\|f(\bm{x};\theta)\bigr{)}$
$\,\leq\rho_{0,\alpha}\left(\theta^{*},\theta\right)$
Thus $S_{0,\alpha}(t)\subseteq S_{p,\alpha}(t)$. ∎
B.3.3 Proof for Lemma 14
Proof.
Note that $m_{n}(A_{n}^{c},\bm{X})=m_{n}(A_{n}^{c}\cap\Omega_{n},\bm{X})+m_{n}(A_{n}^{c}\cap\Omega_{n}^{c},\bm{X})$. From Lemma 9 and 10, we have
$\mathbb{P}^{*}_{\theta^{*}}\left(\sup_{\theta\in A_{n}^{c}\cap\Omega_{n}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta)\geq e^{-c_{1}n\phi_{n}^{2}}\right)\leq C(n,\phi_{n})e^{-c_{2}n\phi_{n}^{2}}.$
Also, by
$m_{n}(A_{n}^{c}\cap\Omega_{n},\bm{X})\leq\sup_{\theta\in A_{n}^{c}\cap\Omega_{n}}\prod_{i=1}^{n}d(\bm{x}_{i},\theta),$
we have
$\mathbb{P}\left(m_{n}(A_{n}^{c}\cap\Omega_{n},\bm{X})\geq e^{-c_{1}n\phi_{n}^{2}}\right)\leq C(n,\phi_{n})e^{-c_{2}n\phi_{n}^{2}}.$
By Markov’s inequality, Fubini’s Theorem and Assumption 3, we have
$\mathbb{P}\left(m_{n}(A_{n}^{c}\cap\Omega_{n}^{c},\bm{X})\geq e^{-c_{1}n\phi_{n}^{2}}\right)\leq$ $\,e^{c_{1}n\phi_{n}^{2}}\iint_{A_{n}^{c}\cap\Omega_{n}^{c}}\prod_{i=1}^{n}d(\bm{x}_{i};\theta)\pi_{0}(\theta)\mathtt{d}\theta\mathtt{d}\mathbb{P}_{\theta^{*}}(\bm{X})$
$=$ $\,e^{c_{1}n\phi_{n}^{2}}\int_{A_{n}^{c}\cap\Omega_{n}^{c}}\int\prod_{i=1}^{n}d(\bm{x}_{i};\theta)\mathtt{d}\mathbb{P}_{\theta^{*}}(\bm{X})\pi_{0}(\theta)\mathtt{d}\theta$
$=$ $\,e^{c_{1}n\phi_{n}^{2}}\mathbb{P}_{\pi_{0}}(A_{n}^{c}\cap\Omega_{n}^{c})\leq c_{4}e^{-c_{1}n\phi_{n}^{2}}.$
Thus
$\mathbb{P}\left(m_{n}(A_{n}^{c},\bm{X})\geq 2e^{-c_{1}n\phi_{n}^{2}}\right)\leq C(n,\phi_{n})e^{-c_{2}n\phi_{n}^{2}}+c_{4}e^{-nc_{1}\phi_{n}^{2}}.$
B.3.4 Proof for Lemma 15
Proof.
Since $f(\bm{x};\theta)$ has all directional derivatives. Let $\phi(\alpha)=\theta+\alpha(\theta^{*}-\theta)$ for some $\alpha\in(0,1)$, note that
$\log k_{p_{n}}(\bm{x};\theta)-\log k_{p_{n}}(\bm{x};\theta^{*})\leq\sup_{\alpha\in(0,1)}\|\tilde{\nabla}_{\theta}\log k_{p_{n}}(\bm{x};\phi(\alpha))\|_{2}\|\theta-\theta^{*}\|_{2}$
where $\tilde{\nabla}$ operator is defined in Definition 7. Note that
~$\|\tilde{\nabla}_{\theta}\log k_{p_{n}}(\bm{x};\theta)\|_{2}=$ $\left\|\frac{(1-p_{n})\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{(1-p_{n})f(\bm{x};\theta)+p_{n}g(\bm{x})}\right\|_{2}$
$\leq$ $\frac{1-p_{n}}{p_{n}}\left\|\frac{\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{g(\bm{x})}\right\|_{2}.$
Since by Assumption 4, ~$\|\tilde{\nabla}_{\theta}f(\bm{x};\theta)/g(\bm{x})\|_{2}$ is bounded as $\|\bm{w}\|_{2}\rightarrow\infty$ in all directions, we have ~$\tilde{\nabla}f(\bm{x};\theta)/g(\bm{x})$ is bounded on $\mathbb{R}^{d}\times A$ for any compact set $A\subset\Omega$. Thus there exists a sequence $T_{n}<\infty$ such that,
$\sup_{x\in\mathbb{R}^{d}}\sup_{\theta\in A_{n}}\left\|\frac{\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{g(\bm{x})}\right\|_{2}\leq T_{n}$
Thus
$|\log d(\bm{x};\theta)|\leq\frac{1-p_{n}}{p_{n}}T_{n}\psi_{n}$
and
$\eta_{n}=\sup_{\bm{x},\bm{z}\in\mathbb{R}^{d}}\sup_{\theta\in A_{n}}\frac{d(\bm{x};\theta)}{d(\bm{z};\theta)}\leq e^{2(1-p_{n})T_{n}\psi_{n}/p_{n}}.$
Appendix C Theoretical Application on Regression Settings
In this section, we will verify the Assumptions in Theorem 5 for linear regression, logistic regression and Cauchy regression. The analysis is presented first, followed by technical lemmas.
C.1 Linear Regression Case
Let
$X_{i}=\theta^{\top}\bm{W}_{i}+e_{i},\qquad i=1,\dots,n$
where $e_{i}$ are assumed to be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables with unknown variance $\sigma^{2}$ and $W_{i}$ are $d$-dimensional covariates with $\|W_{i}\|_{\infty}\leq 1$. The likelihood function of $X_{i}$ given $\theta,\sigma,W_{i}$ can be written as
$f_{\theta,\sigma}(x|\bm{w})=(2\pi\sigma^{2})^{-1/2}\exp\left(-(x-\theta^{\top}\bm{w})^{2}/(2\sigma^{2})\right).$
Firstly, we shall address condition (8) in Assumption 1. To avoid singularity, we choose
$\Omega_{n}:=\{(\theta,\sigma)\in\mathbb{R}^{d}:\|\theta^{*}-\theta\|_{2}\leq R_{n},\,|\sigma-\sigma^{*}|<R_{n},\sigma>1/R_{n}\}$
where $R_{n}\rightarrow\infty$ is defined later.
By Lemma 19, we have for any $(\theta,\sigma),(\vartheta,\varsigma)\in\Omega_{n}$, $\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty}\leq r$,
$\left|\sqrt{f_{\theta,\sigma}(x|\bm{w})}-\sqrt{f_{\vartheta,\varsigma}(x|\bm{w})}\right|\leq M^{(n)}_{\vartheta,\varsigma,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty}$
with
$M^{(n)}_{\vartheta,\varsigma,r}(x,\bm{w})=\frac{R_{n}^{2}}{2}\Bigl{(}d\|\bm{w}\|_{2}+R_{n}(|x|+M_{n}\|\bm{w}\|_{2})\Bigr{)}(|x|+M_{n}\|\bm{w}\|_{2})\\ \left(\frac{R_{n}^{2}}{2\pi}\right)^{1/4}\exp\left(-\frac{1}{4R_{n}^{2}}\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}\right)$
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$. Then by Lemma 21, $\|M^{(n)}_{\vartheta,\varsigma,r}\|^{2}_{2}\preceq M_{n}^{12}$, $\forall(\vartheta,\varsigma)\in\Omega_{n}$ which satisfies the requirement of Lemma 3. Since $\log M_{r}^{(n)}=O(\log M_{n})$, the condition (8) is satisfied by taking
$\phi_{n}=n^{-1/q},\quad M_{n}=n^{k},$
for some $q>4$ and $k>0$ to be chosen for the sake of Assumptions 3.
It is not hard to check that the Gaussian likelihood function satisfies the condition in Lemma 7, when $(\theta,\sigma)$ is close enough to $(\theta^{*},\sigma^{*})$, i.e., within $L_{2}$-distance of some constant $c_{r}$. Thus $S_{0,\alpha}(t_{n})$ contains an $L_{2}$-neighbourhood around $\theta^{*}$ with its radius proportional to $t_{n}$ up to a constant depending on $(\theta^{*},\sigma^{*})$ and $c_{r}$. In addition, Assumption 2 is satisfied by applying Lemma 8 and Remark 14.
Moreover, for $R_{n}$ large, using Remark 15
$\mathbb{P}_{\pi_{0}}(\Omega_{n}^{c})=1-\mathbb{P}_{\pi_{0}}(\Omega_{n})\sim R_{n}^{d-2}e^{-R_{n}^{2}/2}.$
Recall that $M_{n}=R_{n}+\|\theta^{*}\|_{2}$ and $\phi_{n}=n^{-1/q}$, this gives the restriction that
$M_{n}^{2}=n^{2k}\approx R_{n}^{2}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q}$
which can be satisfied by choosing $k>(q-2)/2q$. Finally, Lemma 22 shows that the pair of sequences $\phi_{n}$ and $\psi_{n}$ in Proposition 4 can be chosen (up to a multiplicative constant) to diminish at the same rate when $(\theta,\sigma)$ is close to $(\theta^{*},\sigma^{*})$. Therefore $\psi_{n}$ can also be chosen to be $\propto\phi_{n}=n^{-1/q}$. By choosing $p_{n}$ to converge at a rate slower than $n^{-1/q}$, e.g., $p_{n}=n^{-1/2q}$, then
$\epsilon_{n}=O(n^{-1/2q}),\qquad\delta_{n}=O(\exp(-\tfrac{1}{2}c_{1}n^{1-2/q}))$
where $c_{1}$ is a fixed constant and $q>4$.
Lemma 19.
Let
$f_{\theta,\sigma}(x|\bm{w}):=(2\pi\sigma^{2})^{-1/2}\exp(-(x-\theta^{\top}\bm{w})^{2}/(2\sigma^{2})),$
where $(\theta,\sigma)\in\Omega_{n}:=\{(\theta,\sigma)\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n},|\sigma^{*}-\sigma|\leq R_{n},\sigma>1/R_{n}\}$, for some $R_{n}>0$. If $(\theta,\sigma),(\vartheta,\varsigma)\in\Omega_{n}$ with $\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty}\leq r$ then
$\left|\sqrt{f_{\theta,\sigma}(x|\bm{w})}-\sqrt{f_{\vartheta,\varsigma}(x|\bm{w})}\right|\leq M_{\vartheta,\varsigma,r}(x,\bm{w})\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty}$
where function $M_{\vartheta,\varsigma,r}$ is given by
$M_{\vartheta,\varsigma,r}(x,\bm{w}):=\frac{R_{n}^{2}}{2}\Bigl{(}d\|\bm{w}\|_{2}+R_{n}(|x|+M_{n}\|\bm{w}\|_{2})\Bigr{)}(|x|+M_{n}\|\bm{w}\|_{2})\\ \left(\frac{R_{n}^{2}}{2\pi}\right)^{1/4}\exp\left(-\frac{1}{4R_{n}^{2}}\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}\right)$
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$.
Proof.
Recall that
$f_{\theta,\sigma}(x|\bm{w})=(2\pi\sigma^{2})^{-1/2}\exp(-(x-\theta^{\top}\bm{w})^{2}/(2\sigma^{2})),$
Then, by Cauchy’s Mean Value Theorem and Cauchy-Schwarz inequality, there exists some $a\in(0,1)$, $\varphi=\vartheta+a(\theta-\vartheta)$, $\zeta=\sigma_{2}+a(\sigma_{2}-\sigma_{1})$, such that
$\,\left|\sqrt{f_{\theta,\sigma}(x|\bm{w})}-\sqrt{f_{\vartheta,\varsigma}(x|\bm{w})}\right|$
$=$ $\,\frac{x-\varphi^{\top}\bm{w}}{2\zeta^{2}}f_{\varphi,\zeta}^{\frac{1}{2}}(x|\bm{w})\,(\theta-\vartheta)^{\top}\bm{x}+\frac{1}{2}\left(\frac{(x-\varphi^{\top}\bm{w})^{2}}{\zeta^{3}}-\zeta^{-1}\right)f_{\varphi,\zeta}^{\frac{1}{2}}(x|\bm{w})\,(\sigma_{1}-\sigma_{2})$
$\leq$ $\,\frac{1}{2\zeta^{2}}\left((x-\varphi^{\top}\bm{w})d\|\bm{w}\|_{2}+\frac{(x-\varphi^{\top}\bm{w})^{2}}{\zeta}\right)f_{\varphi,\zeta}^{\frac{1}{2}}(x|\bm{w})\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty}$
$\leq$ $\,\frac{R_{n}^{2}}{2}\Bigl{(}d\|\bm{w}\|_{2}+R_{n}(|x|+M_{n}\|\bm{w}\|_{2})\Bigr{)}(|x|+M_{n}\|\bm{w}\|_{2})$
×$\,\times\left(\frac{R_{n}^{2}}{2\pi}\right)^{1/4}\exp\left(-\frac{1}{4R_{n}^{2}}\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}\right)\|(\theta,\sigma),(\vartheta,\varsigma)\|_{\infty}.$
The last step is by noting the following upper bounds since $(\varphi,\zeta)\in\Omega_{n}$
$|x-\varphi^{\top}\bm{w}|\leq|x|+M_{n}\|\bm{w}\|_{2},\quad\frac{1}{\zeta}\leq R_{n}.$
Lemma 20.
For constants $R,\alpha\in(0,M]$, $k\in\mathbb{Z}_{+}$,
$\int_{0}^{\infty}x^{k}\frac{1}{\sqrt{2\pi R^{2}}}\exp\left(-\frac{(x-\alpha w)^{2}}{2R^{2}}\right)\mathtt{d}x\leq\beta_{k}(w)M^{k}$
where
$\beta_{k}(w)=\frac{1}{\sqrt{\pi}}\left[2\sum_{j=0}^{\left\lfloor\frac{k}{2}\right\rfloor}\begin{pmatrix}k\\ 2j\end{pmatrix}2^{j-1}w^{k-2j}\Gamma(\tfrac{2j+1}{2})+\sum_{j=1}^{\left\lceil\frac{k}{2}\right\rceil}\begin{pmatrix}k\\ 2j-1\end{pmatrix}2^{\frac{2j-1}{2}-1}w^{k-2j+1}\Gamma(j)\right].$
Proof.
First, we compute the following integral
$\,\int_{0}^{\infty}\frac{w^{k}}{\sqrt{2\pi R^{2}}}\exp\left(-\frac{w^{2}}{2R^{2}}\right)\mathtt{d}w$
$=$ $\,\frac{1}{\sqrt{2\pi}}\int_{0}^{\infty}(2R^{2}t)^{\frac{k-1}{2}}R\exp(-t)\mathtt{d}t$ $[t=\tfrac{w^{2}}{2R^{2}}]$
$=$ $\,\frac{1}{\sqrt{\pi}}2^{k/2-1}R^{k}\Gamma(\tfrac{k+1}{2}).$
Therefore,
$\,\int_{0}^{\infty}x^{k}\frac{1}{\sqrt{2\pi R^{2}}}\exp\left(-\frac{(x-\alpha w)^{2}}{2R^{2}}\right)\mathtt{d}x$
$=$ $\,\int_{-\alpha w}^{\infty}(t+\alpha w)^{k}\frac{1}{\sqrt{2\pi R^{2}}}\exp\left(-\frac{t^{2}}{2R^{2}}\right)\mathtt{d}t$ $[t=x-\alpha w]$
$\leq$  $\,2\sum_{j=0}^{\left\lfloor\frac{k}{2}\right\rfloor}\begin{pmatrix}k\\ 2j\end{pmatrix}\int_{0}^{\infty}t^{2j}(\alpha w)^{k-2j}\frac{1}{\sqrt{2\pi R^{2}}}\exp\left(-\frac{t^{2}}{2R^{2}}\right)\mathtt{d}t$
$\,+\,\sum_{j=1}^{\left\lceil\frac{k}{2}\right\rceil}\begin{pmatrix}k\\ 2j-1\end{pmatrix}\int_{0}^{\infty}t^{2j-1}(\alpha w)^{k-2j+1}\frac{1}{\sqrt{2\pi R^{2}}}\exp\left(-\frac{t^{2}}{2R^{2}}\right)\mathtt{d}t$
$\leq$ $\,\beta_{k}(w)M^{k}$
where
$\beta_{k}(w)=\frac{1}{\sqrt{\pi}}\left[2\sum_{j=0}^{\left\lfloor\frac{k}{2}\right\rfloor}\begin{pmatrix}k\\ 2j\end{pmatrix}2^{j-1}w^{k-2j}\Gamma(\tfrac{2j+1}{2})+\sum_{j=1}^{\left\lceil\frac{k}{2}\right\rceil}\begin{pmatrix}k\\ 2j-1\end{pmatrix}2^{\frac{2j-1}{2}-1}w^{k-2j+1}\Gamma(j)\right].$
Lemma 21.
Define
$\Omega_{n}:=\{(\theta,\sigma)\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n},|\sigma^{*}-\sigma|\leq R_{n},\sigma>1/R_{n}\},$
for some $R_{n}>0$. For $(\vartheta,\varsigma)\in\Omega_{n}$, let
$M_{\vartheta,\varsigma,r}(x,\bm{w}):=\frac{R_{n}^{2}}{2}\Bigl{(}d\|\bm{w}\|_{2}+R_{n}(|x|+M_{n}\|\bm{w}\|_{2})\Bigr{)}(|x|+M_{n}\|\bm{w}\|_{2})\\ \left(\frac{R_{n}^{2}}{2\pi}\right)^{1/4}\exp\left(-\frac{1}{4R_{n}^{2}}\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}\right),$
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$. If $\|\bm{w}\|_{\infty}\leq 1$, then
$\|M_{\vartheta,\varsigma,r}(x,\bm{w})\|_{2}^{2}\preceq M_{n}^{12}.$
Proof.
$\|M_{\vartheta,\varsigma,r}(x,\bm{w})\|_{2}^{2}=\int M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}x.$
Consider the value of
$\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w}).$
Note that $\varphi=\vartheta+\tilde{r}\bm{v}$ for some $\tilde{r}<r$ and unit vector $\bm{v}$. When $\bm{w}$ is fixed, $\varphi^{\top}\bm{w}$ takes value within the interval $(u_{-},u_{+})$ where
$u_{-}:=\vartheta^{\top}\bm{w}-r\|\bm{w}\|_{2},\quad u_{+}:=\vartheta^{\top}\bm{w}+r\|\bm{w}\|_{2},$
with the infimum and the supremum attained at $\vartheta-\frac{r\bm{w}}{\|\bm{w}\|_{2}}$ and $\vartheta+\frac{r\bm{w}}{\|\bm{w}\|_{2}}$ respectively. Then
$\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}=\begin{dcases}(x-\vartheta^{\top}\bm{w}+r\|\bm{w}\|_{2})^{2},&x<u_{-};\\ 0,&u_{-}\leq x\leq u_{+};\\ (x-\vartheta^{\top}\bm{w}-r\|\bm{w}\|_{2})^{2},&x>u_{+};\end{dcases}$
Write
$\int_{-\infty}^{\infty}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}x=\int_{-\infty}^{u_{-}}+\int_{u_{-}}^{u_{+}}+\int_{u_{+}}^{\infty}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}x.$
Firstly, by noting that $|u_{+}|\leq M_{n}\|\bm{w}\|_{2}$
$\,\int_{u_{-}}^{u_{+}}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}x$
$=$ $\,\frac{1}{\sqrt{2\pi}}\frac{R_{n}^{5}}{4}\int_{u-}^{u_{+}}\Bigl{(}d\|\bm{w}\|_{2}+R_{n}(|x|+M_{n}\|\bm{w}\|_{2})\Bigr{)}^{2}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\exp\left(-\frac{1}{2R_{n}^{2}}\right)\mathtt{d}x$
$\leq$ $\,\frac{1}{\sqrt{2\pi}}\frac{M_{n}^{5}}{4}(d+2M_{n}^{2})^{2}\|\bm{w}\|_{2}^{2}(2M_{n}\|\bm{w}\|_{2})^{2}2r\|\bm{w}\|_{2}$
$\preceq$ $M_{n}^{11}.$
Note that $\int_{-\infty}^{u_{-}}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}y$ is comprised of summing integrals of form
$\int_{-\infty}^{u_{-}}\frac{1}{\sqrt{2\pi R_{n}^{2}}}|x|^{k}\exp\left(-\frac{(x-u_{-})^{2}}{2R_{n}^{2}}\right)\mathtt{d}x.$
By Lemma 20,
$\,\int_{-\infty}^{u_{-}}\frac{1}{\sqrt{2\pi R_{n}^{2}}}|x|^{k}\exp\left(-\frac{(x-u_{-})^{2}}{2R_{n}^{2}}\right)\mathtt{d}x$
$\leq$ $\,\int_{-\infty}^{\infty}\frac{1}{\sqrt{2\pi R_{n}^{2}}}|x|^{k}\exp\left(-\frac{(x-u_{-})^{2}}{2R_{n}^{2}}\right)\mathtt{d}x$
$\leq$  $\,2\int_{0}^{\infty}\frac{1}{\sqrt{2\pi R_{n}^{2}}}x^{k}\exp\left(-\frac{(x-|u_{-}|)^{2}}{2R_{n}^{2}}\right)\mathtt{d}x$
$\preceq$  $\,2\beta_{k}(\|\bm{w}\|_{2})M_{n}^{k}.$
Since $|u_{-}|,|u_{+}|\leq M_{n}\|\bm{w}\|_{2}$, the integrals $\int_{-\infty}^{u_{-}}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}y$ and $\int_{u_{-}}^{\infty}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}y$ share the same upper bound as above (see by change of variable $w=-y$). Then $\|M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\|_{2}^{2}\preceq M_{n}^{12}$.
Lemma 22.
Let $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
$f_{\theta,\sigma}(x|\bm{w}):=(2\pi\sigma^{2})^{-1/2}\exp\left(-\frac{(x-\theta^{\top}\bm{w})^{2}}{2\sigma^{2}}\right).$
Then we have for $(\theta,\sigma)$ close to $(\theta^{*},\sigma^{*})$, there exists positive constant $C$ independent of $(\theta,\sigma)$ such that
$h((\theta^{*},\sigma^{*}),(\theta,\sigma))\leq\phi_{n}\implies\|(\theta^{*},\sigma^{*})-(\theta,\sigma))\|_{2}\leq C\phi_{n}$
Proof.
Recall that by Cauchy’s mean value theorem, there exists $\varphi=\theta^{*}+\alpha(\theta-\theta^{*})$, $\xi=\sigma^{*}+\alpha(\sigma-\sigma^{*})$, $\alpha\in(0,1)$ such that
$\,f^{\frac{1}{2}}_{\theta,\sigma}(x|\bm{w})-f^{\frac{1}{2}}_{\theta^{*},\sigma^{*}}(x|\bm{w})$
$=$ $\,f_{\varphi,\xi}^{\frac{1}{2}}(x|\bm{w})\left[\frac{x-\varphi^{\top}\bm{w}}{2\xi^{2}}(\theta-\theta^{*})^{\top}\bm{x}+\frac{1}{2}\left(\frac{(x-\varphi^{\top}\bm{w})^{2}}{\xi^{3}}-\xi^{-1}\right)(\sigma-\sigma^{*})\right].$
Let $z:=(x-\varphi^{\top}\bm{w})/\xi$, then
$h^{2}((\theta^{*},\sigma^{*}),(\theta,\sigma)):=$ $\,\frac{1}{2}\int\left(f^{\frac{1}{2}}_{\theta,\sigma}(x|\bm{w})-f^{\frac{1}{2}}_{\theta^{*},\sigma^{*}}(x|\bm{w})\right)^{2}\mathtt{d}x$
$=$ $\,\frac{1}{8\xi^{2}}\int f_{\varphi,\xi}(x|\bm{w})\left[z(\theta-\theta^{*})^{\top}\bm{w}+\left(z^{2}-1\right)(\sigma-\sigma^{*})\right]^{2}\mathtt{d}x.$
Note that in the above integral with respect to $x$, $z$ is a standard normal random variable and $f_{\varphi,\xi}(x|\bm{w})\propto f_{z}(z)$. There exists constant $\lambda$ depending only on the (maximum) distance between $(\theta,\sigma)$ and $(\theta^{*},\sigma^{*})$ such that
$h^{2}((\theta^{*},\sigma^{*}),(\theta,\sigma))\geq\lambda\|(\theta,\sigma)-(\theta^{*},\sigma^{*})\|_{2}^{2}.$
C.2 Logistic Regression Case
Consider the following linear logistic regression model,
$X_{i}\sim\text{Bern}(p_{i}),\qquad p_{i}=\Phi(\theta^{\top}\bm{W}_{i}),\quad\theta,\bm{W}_{i}\in\mathbb{R}^{d},$
where $\Phi(z)=(1+e^{-z})^{-1}$. Again, we assume the covariates satisfy $\|\bm{W}_{i}\|_{\infty}\leq 1$, then
$f_{\theta}(x|\bm{w})=\Phi(\theta^{\top}\bm{w})^{x}\Phi(-\theta^{\top}\bm{w})^{1-x}.$
Note that
$\left|\sqrt{f_{\theta}(x|\bm{w})}-\sqrt{f_{\vartheta}(x|\bm{w})}\right|\leq\frac{1}{2}\|\bm{w}\|_{2}\|\theta-\vartheta\|_{2}.$
So we can apply Lemma 3 with $M^{(n)}_{\vartheta,r}(x,\bm{w})=\frac{1}{2}\|\bm{w}\|_{2}$. Hence $\|M^{(n)}_{\vartheta,r}(x,\bm{w})\|_{2}^{2}=O(d)$ is bounded above by a uniformed constant since $d$ is fixed. Thus, Assumption 1 can be satisfied with
$\phi_{n}=n^{-1/q},\quad R_{n}=n^{k}$
for some $q>4$ and $k>0$ to be chosen for Assumption 3. Again, it is not hard to check Lemma 7 holds for this likelihood and thus Assumption 2 is satisfied by applying Lemma 8 and Remark 14. Moreover, for $R_{n}$ large, by Remark 15,
$\mathbb{P}_{\pi_{0}}(\Omega_{n}^{c})=1-\mathbb{P}_{\pi_{0}}(\Omega_{n})\sim R_{n}^{d-2}e^{-R_{n}^{2}/2}.$
Recall that $M_{n}=R_{n}+\|\theta^{*}\|_{2}$ and $\phi_{n}=n^{-1/q}$, this gives the restriction again that
$M_{n}^{2}=n^{2k}\approx R_{n}^{2}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q}$
which can be satisfied by choosing $k>(q-2)/2q$ similar to the previous example.
Finally, Lemma 23 shows that the pair of sequences $\phi_{n}$ and $\psi_{n}$ in Proposition 4 can be chosen (up to a multiplicative constant) to diminish at the same rate when $\theta$ is close to $\theta^{*}$. Therefore $\psi_{n}$ can also be chosen to be $\propto\phi_{n}=n^{-1/q}$. By choosing $p_{n}$ to converge at a rate slower than $n^{-1/q}$, e.g., $p_{n}=n^{-1/2q}$, then
$\epsilon_{n}=O(n^{-1/2q}),\qquad\delta_{n}=O(\exp(-\tfrac{1}{2}c_{1}n^{1-2/q}))$
where $c_{1}$ is a fixed constant and $q>4$.
Lemma 23.
Let $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
$f_{\theta}(x|\bm{w}):=\Phi(\theta^{\top}\bm{w})^{x}\Phi(-\theta^{\top}\bm{w})^{1-x}$
where $\Phi(x)=(1+e^{-x})^{-1}$. Then we have for $\theta$ close to $\theta^{*}$, there exists positive constant $C$ independent of $\theta$ such that
$h(\theta^{*},\theta)\leq\phi_{n}\implies\|\theta^{*}-\theta\|_{2}\leq C\phi_{n}$
Proof.
Note that
$\nabla_{\theta}f^{\frac{1}{2}}_{\theta}(x|\bm{w})=\left(\frac{x}{2}\Phi(\theta^{\top}\bm{w})^{\frac{x}{2}}\Phi(-\theta^{\top}\bm{w})^{\frac{3-x}{2}}-\frac{1-x}{2}\Phi(\theta^{\top}\bm{w})^{\frac{x+2}{2}}\Phi(-\theta^{\top}\bm{w})^{\frac{1-x}{2}}\right)\bm{w}.$
Then there exists $\varphi=\theta^{*}+\alpha(\theta-\theta^{*})$, $\alpha\in(0,1)$ such that
$\sum_{y=0}^{1}\left(f^{\frac{1}{2}}_{\theta^{*}}(x|\bm{w})-f^{\frac{1}{2}}_{\theta}(x|\bm{w})\right)^{2}=\frac{1}{4}\Phi(\varphi^{\top}\bm{w})\Phi(-\varphi^{\top}\bm{w})\left((\theta-\theta^{*})^{\top}\bm{w}\right)^{2}.$
Now, by Assumption $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries
$h^{2}(\theta^{*},\theta)=\frac{1}{8}\Phi(\varphi^{\top}\bm{w})\Phi(-\varphi^{\top}\bm{w})\left((\theta-\theta^{*})^{\top}\bm{w}\right)^{2}$
Since $\theta$ is assumed to be close to $\theta^{*}$, $\varphi$ takes value in a compact set, and $\bm{w}\|_{\infty}\leq 1$, hence $\exists C>0$ such that
$h(\theta^{*},\theta)\geq C\|\theta^{*}-\theta\|_{2}.$
C.3 Cauchy Regression Case
Consider the following Cauchy regression model,
$X_{i}=\theta^{\top}\bm{W}_{i}+e_{i},$
where $e_{i}\sim\text{Cauchy}(0,\gamma)$ is Cauchy random variable with unknown dispersion parameter $\gamma$. Assume that the covariates $\bm{W}_{i}$ satisfy $\|\bm{W}_{i}\|_{\infty}\leq 1$. Then the likelihood is given by
$f_{\theta,\gamma}(x|\bm{w})=\frac{\sqrt{\gamma}}{\pi}\left(1+\frac{(x-\theta^{\top}\bm{w})^{2}}{\gamma}\right)^{-1}.$
By Cauchy’s mean value theorem, we can show (as in Lemma 24 that for any $(\theta,\gamma_{1}),(\vartheta,\gamma_{2})\in\Omega_{n}:=\{(\vartheta,\gamma)\in\Omega:\|\vartheta-\theta^{*}\|_{2}\leq R_{n},|\gamma-\gamma^{*}|<R_{n},\gamma>R_{n}^{-1}]\}$, such that $\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty}\leq r$,
$\left|\sqrt{f_{\theta,\gamma_{1}}(x|\bm{w})}-\sqrt{f_{\vartheta,\gamma_{2}}(x|\bm{w})}\right|\leq M^{(n)}_{\vartheta,\gamma_{2},r}(x,\bm{w})\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty}$
where
$M^{(n)}_{\vartheta,\gamma_{2},r}(x,\bm{w})=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]$
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$, $B_{r}(\vartheta)$ is the $L_{2}$-ball around $\vartheta$ and by Lemma 25, $\|M_{\vartheta,r}\|_{2}^{2}\preceq M_{n}^{9}$. Thus, similar to the first example, $\log M_{r}^{(n)}=O(\log M_{n})$ and Assumption 1 can be satisfied by taking
$\phi_{n}=n^{-1/q},\quad M_{n}=n^{k}$
for some $q>4$ and $k>0$ to be chosen for Assumption 3. It is not hard to check Lemma 7 holds for this likelihood and thus Assumption 2 is satisfied by applying Lemma 8 and Remark 14. Assumption 3 can be checked using the same procedure as in the Linear regression example which again gives $k>\tfrac{1}{4}$.
Finally, Lemma 26 shows that the pair of sequences $\phi_{n}$ and $\psi_{n}$ in Proposition 4 can be chosen (up to a multiplicative constant) to diminish at the same rate when $\theta$ is close to $\theta^{*}$. Therefore $\psi_{n}$ can also be chosen to be $\propto\phi_{n}=n^{-1/q}$. By choosing $p_{n}$ to converge at a rate slower than $n^{-1/q}$, e.g., $p_{n}=n^{-1/2q}$, then
$\epsilon_{n}\sim n^{-1/2q},\qquad\delta_{n}\sim\exp(-\tfrac{1}{2}c_{1}n^{1-2/q})$
where $c_{1}$ is a fixed constant, and $q>4$.
Lemma 24.
Let
$f_{\theta,\gamma}(x|\bm{w}):=\frac{\sqrt{\gamma}}{\pi}(1+(x-\theta^{\top}\bm{w})^{2}/\gamma)^{-1},$
where $(\theta,\gamma)\in\Omega_{n}:=\{(\theta,\gamma)\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n},|\gamma-\gamma^{*}|<R_{n},\gamma>R_{n}^{-1}]\}$, $R_{n}>0$. If $\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty}\leq r$ then
$\left|\sqrt{f_{\theta,\gamma_{1}}(x|\bm{w})}-\sqrt{f_{\vartheta,\gamma_{2}}(x|\bm{w})}\right|\leq M_{\vartheta,\gamma_{2},r}(x,\bm{w})\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty}$
where function $M_{\vartheta,\gamma_{2},r}$ is given by
$M_{\vartheta,\gamma_{2},r}(x,\bm{w}):=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]$
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$, $B_{r}(\vartheta)$ is the $L_{2}$-ball around $\vartheta$.
Proof.
$\nabla_{\theta}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\gamma^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\frac{x-\theta^{\top}\bm{w}}{\gamma}\bm{w}$
$\nabla_{\gamma}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\frac{1}{4}\gamma^{-\frac{7}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\left[\gamma+3(x-\theta^{\top}\bm{w})^{2}\right]$
By Cauchy’s mean value theorem, there exists $\varphi=\vartheta+\alpha(\theta-\vartheta)$, $\xi=\gamma_{2}+\alpha(\gamma_{1}-\gamma_{2})$ for some $\alpha\in(0,1)$ such that
$\,\left|\sqrt{f_{\theta,\gamma_{1}}(x|\bm{w})}-\sqrt{f_{\vartheta,\gamma_{2}}(x|\bm{w})}\right|$
$=$ $\,\xi^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\varphi^{\top}\bm{w})^{2}/\xi\right)^{-\frac{3}{2}}\frac{x-\varphi^{\top}\bm{w}}{\xi}(\theta-\vartheta)^{\top}\bm{x}$
$\quad+\frac{1}{4}\xi^{-\frac{7}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\varphi^{\top}\bm{w})^{2}/\xi\right)^{-\frac{3}{2}}\left[\xi+3(x-\varphi^{\top}\bm{w})^{2}\right](\gamma_{1}-\gamma_{2})$
$\leq$ $\,\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\cdot\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty}$
$\quad\cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]$
Lemma 25.
For $(\theta,\gamma)\in\Omega_{n}:=\{(\theta,\gamma)\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n},\gamma\in[R_{n}^{-1},R_{n}]\}$, $R_{n}>0$, let
$M_{\vartheta,\gamma_{2},r}(x,\bm{w}):=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]$
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$, $B_{r}(\vartheta)$ is the $L_{2}$-ball around $\vartheta$. If $\|\bm{w}\|_{\infty}\leq 1$, then
$\|M_{\vartheta,r}(x,\bm{w})\|_{2}^{2}\preceq M_{n}^{9}.$
Proof.
Similar to Lemma 21, we consider the value of $\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}$ in different configurations of $(x,\bm{w},\vartheta)$. Since $\varphi=\vartheta+\tilde{r}v$ for some $\tilde{r}<r$ and some unit vector $v$, when $x$ is fixed, $\varphi^{\top}\bm{w}$ takes value within the interval $(u_{-},u_{+})$ where
$u_{-}:=\vartheta^{\top}\bm{w}-r\|\bm{w}\|_{2},\qquad u_{+}:=\vartheta^{\top}\bm{w}+r\|\bm{w}\|_{2},$
with the infimum and supremum attained at $\vartheta-\frac{r\bm{w}}{\|\bm{w}\|_{2}}$ and $\vartheta+\frac{r\bm{w}}{\|\bm{w}\|_{2}}$ respectively. Therefore $(x-\varphi^{\top}\bm{w})^{2}$ is minimized by choosing $\varphi=\vartheta-r\bm{w}$ when $x<u_{-}$ and $\varphi=\vartheta+r\bm{w}$ when $x>u_{+}$. When $x\in(u_{-},u_{+})$, $(x-\varphi^{\top}\bm{w})^{2}$ has minimum zero. For simplicity, let
$S(x,\bm{w})=\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]^{2}.$
Again, the integral $\|M_{\vartheta,\gamma_{2},r}(x,\bm{w})\|_{2}^{2}$ can be evaluated separately in three intervals $\int_{-\infty}^{u_{-}}$,$\int_{u_{-}}^{u_{+}}$ and $\int_{u_{+}}^{\infty}$. By noting that $\max\{|u_{-}|,|u_{+}|\}\leq(\|\vartheta\|_{2}+r)\|\bm{w}\|_{2}\leq M_{n}\|\bm{w}\|_{2}$,
$\int_{u_{-}}^{u_{+}}\frac{1}{\pi}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-3}S(x,\bm{w})\mathtt{d}x\preceq M_{n}^{\frac{17}{2}}\|\bm{w}\|_{2}^{5}\preceq M_{n}^{\frac{17}{2}}.$
Moreover, since the integrand is non-negative, for $k\leq 4$,
$\,\int_{-\infty}^{u_{-}}\frac{1}{\pi}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-3}|x|^{k}\mathtt{d}x$
$\leq$ $\,\int_{-\infty}^{\infty}\frac{1}{\pi}\left(1+(x-u_{-})^{2}/R_{n}\right)^{-3}|x|^{k}\mathtt{d}x$
$=$ $\,\int_{0}^{\infty}\frac{1}{\pi}\left(1+(x-u_{-})^{2}/R_{n}\right)^{-3}x^{k}\mathtt{d}x+\int_{0}^{\infty}\frac{1}{\pi}\left(1+(x+u_{-})^{2}/R_{n}\right)^{-3}x^{k}\mathtt{d}x$
$\preceq$ $\,M_{n}^{k+\frac{1}{2}}.$
The last step is by computing the integral
$\,\int_{0}^{\infty}\frac{1}{\pi}\left(1+(x-a)^{2}/R_{n}\right)^{-3}x^{k}\mathtt{d}x$
$=$ $\,\int_{0}^{\infty}\frac{1}{\pi}\left(1+t^{2}/R_{n}\right)^{-3}(t+a)^{k}\mathtt{d}t$ $[t=y-a]$
$\leq$  $\,2\sum_{j=0}^{\left\lfloor\frac{k}{2}\right\rfloor}\binom{k}{2j}\int_{0}^{\infty}t^{2j}a^{k-2j}\frac{1}{\pi}(1+t^{2}/R_{n})^{-3}\mathtt{d}t$
$\quad+\sum_{j=1}^{\left\lceil\frac{k}{2}\right\rceil}\binom{k}{2j-1}\int_{0}^{\infty}t^{2j-1}a^{k-2j-1}\frac{1}{\pi}(1+t^{2}/R_{n})^{-3}\mathtt{d}t$
$\preceq$ $\,\max\{a,R_{n}\}^{k+\frac{1}{2}}$
while noting
$\int_{0}^{\infty}\frac{1}{\pi}t^{k}(1+t^{2}/R_{n})^{-3}\mathtt{d}t\leq R_{n}^{\frac{k+1}{2}}.$
Therefore,
$\int_{-\infty}^{u_{-}}\frac{1}{\pi}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-3}S(x,\bm{w})\mathtt{d}x\preceq M_{n}^{9}.$
The same upper bound applies to
$\int_{u_{+}}^{\infty}\frac{1}{\pi}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-3}S(x,\bm{w})\mathtt{d}x.$
Hence,
$\|M_{\vartheta,\gamma_{2},r}(x,\bm{w})\|_{2}^{2}\preceq M_{n}^{9}.$
Lemma 26.
Let $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
$f_{\theta,\gamma}(x|\bm{w}):=\frac{\sqrt{\gamma}}{\pi}(1+(x-\theta^{\top}\bm{w})^{2}/\gamma)^{-1}.$
Then we have for $(\theta,\gamma)$ close to $(\theta^{*},\gamma^{*})$, there exists positive constant $C$ independent of $(\theta,\gamma)$ such that
$h((\theta^{*},\gamma^{*}),(\theta,\gamma))\leq\phi_{n}\implies\|(\theta^{*},\gamma^{*})-(\theta,\gamma))\|_{2}\leq C\phi_{n}$
Proof.
$h^{2}((\theta^{*},\gamma^{*}),(\theta,\gamma))=\frac{1}{2}\int\left(\sqrt{\frac{\sqrt{\gamma^{*}}}{\pi(1+(x-\theta^{*\top}\bm{w})^{2}/\gamma^{*})}}-\sqrt{\frac{\sqrt{\gamma}}{\pi(1+(x-\theta^{\top}\bm{w})^{2}/\gamma)}}\right)^{2}\mathtt{d}y.$
Recall the grad of $\sqrt{f_{\theta,\gamma}}$ computed in Lemma 24.
$\nabla_{\theta}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\gamma^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\frac{x-\theta^{\top}\bm{w}}{\gamma}\bm{w}$
$\nabla_{\gamma}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\frac{1}{4}\gamma^{-\frac{7}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\left[\gamma+3(x-\theta^{\top}\bm{w})^{2}\right]$
By Cauchy’s mean value theorem, there exists $\varphi=\theta^{*}+\alpha(\theta-\theta^{*})$, $\xi=\gamma^{*}+\alpha(\gamma-\gamma^{*})$ for some $\alpha\in(0,1)$ such that
$\,\int\left(\sqrt{\frac{\sqrt{\gamma^{*}}}{\pi(1+(x-\theta^{*\top}\bm{w})^{2}/\gamma^{*})}}-\sqrt{\frac{\sqrt{\gamma}}{\pi(1+(x-\theta^{\top}\bm{w})^{2}/\gamma)}}\right)^{2}\mathtt{d}x$
$=$ $\,\int\frac{\sqrt{\xi}}{\pi}\left(1+\frac{(x-\varphi^{\top}\bm{w})^{2}}{\xi}\right)^{-3}\left[\frac{x-\varphi^{\top}\bm{w}}{\xi}(\theta-\theta^{*})^{\top}\bm{x}+\frac{\gamma-\gamma^{*}}{4\xi}+\frac{3}{\xi^{2}}(x-\varphi^{\top}\bm{w})^{2}(\gamma-\gamma^{*})\right]^{2}\mathtt{d}x$
$\succeq$ $\,\left((\theta-\theta^{*})^{\top}\bm{w}\right)^{2}+(\gamma-\gamma^{*})^{2}.$
The last step is due to the value of integral with respect to $\frac{x-\varphi^{\top}\bm{w}}{\xi}$ is only dependent on $\xi$ which is ‘close’ to $\gamma^{*}$ and thus has a finite lower and upper bound. Thus, since $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
$h^{2}((\theta^{*},\gamma^{*}),(\theta,\gamma))\succeq\|(\theta,\gamma)-(\theta^{*}-\gamma^{*})\|_{2}^{2}.$
C.4 Quantile Regression Case
Lemma 27.
Let
$f^{\tau}_{\theta}(x|\bm{w}):=\tau(1-\tau)\exp\left(-Q_{\tau}(x-\theta^{\top}\bm{w})\right),$
where $\theta\in\Omega_{n}:=\{\theta\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n}\}$, for some $R_{n}>0$,
$Q_{\tau}(u):=\begin{dcases}(\tau-1)u,&u\leq 0\\ \tau u,&u>0.\end{dcases}$
If $\theta,\vartheta\in\Omega_{n}$ with $\|\theta-\vartheta\|_{\infty}\leq r$ then
$\left|\sqrt{f^{\tau}_{\theta}(x|\bm{w})}-\sqrt{f^{\tau}_{\vartheta}(x|\bm{w})}\right|\leq M_{\vartheta,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty}$
where function $M_{\vartheta,r}$ is given by
$M_{\vartheta,r}(x,\bm{w}):=\frac{1}{2}\exp\left(-\frac{1-\tau^{*}}{2}\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|\right)\|\bm{w}\|_{2}.$
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$, $\tau^{*}=\max\{\tau,1-\tau\}$.
Proof.
Note that
$\nabla_{\theta}\sqrt{f^{\tau}_{\theta}(x|\bm{w})}=\begin{dcases}-\sqrt{\tau(1-\tau)}(1-\tau)\exp\left(-\frac{\tau-1}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&y\leq\theta^{\top}\bm{w}\\ -\sqrt{\tau(1-\tau)}\tau\exp\left(-\frac{\tau}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&y>\theta^{\top}\bm{w}.\end{dcases}$
Using the same argument as in Lemma 20, recall the $\tilde{\nabla}$ operator from Definition 9 for the supremum of all direction derivatives. Let $\varphi=\theta+\alpha(\vartheta-\theta)$, then
$\sqrt{f^{\tau}_{\theta}(x|\bm{w})}-\sqrt{f^{\tau}_{\vartheta}(x|\bm{w})}\leq$ $\,\sup_{\alpha\in(0,1)}\left\|\tilde{\nabla}_{\theta}\sqrt{f_{\varphi}^{\tau}(x|\bm{w})}\right\|_{2}\|\theta-\vartheta\|_{2}$
$\leq$ $\,\sup_{\varphi\in B_{r}(\vartheta)}\frac{d\tau^{*}}{2}\exp\left(-\frac{1-\tau^{*}}{2}\left|x-\varphi^{\top}\bm{w}\right|\right)\|\bm{w}\|_{2}\|\theta-\vartheta\|_{\infty}.$
Thus, we can take
$M_{\vartheta,r}(x,\bm{w})=\frac{d}{2}\exp\left(-\frac{1-\tau^{*}}{2}\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|\right)\|\bm{w}\|_{2}.$
Lemma 28.
Define
$\Omega_{n}:=\{\theta\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n}\}$
for some $R_{n}>0$. For $(\vartheta,\varsigma)\in\Omega_{n}$, let
$M_{\vartheta,r}(x,\bm{w}):=\frac{d}{2}\exp\left(-\frac{1-\tau^{*}}{2}\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|\right)\|\bm{w}\|_{2},$
where $\tau^{*}=\max\{\tau,1-\tau\}$. If $\|\bm{w}\|_{\infty}\leq 1$, then
$\|M_{\vartheta,r}(x,\bm{w})\|_{2}^{2}=\frac{d^{2}}{2(1-\tau^{*})}\|\bm{w}\|_{2}^{2}+\frac{d^{2}r}{2}\|\bm{w}\|_{2}^{3}=O(1).$
Proof.
$\|M_{\vartheta,r}\|_{2}^{2}=\frac{d^{2}}{4}\|\bm{w}\|_{2}^{2}\int\exp\left(-(1-\tau^{*})\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|\right)\mathtt{d}x.$
The integral can be treated similarly to the Gaussian or the Cauchy case by considering the value of $\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|$ in different cases. Define
$u_{-}:=\vartheta^{\top}\bm{w}-r\|\bm{w}\|_{2},\quad u_{+}:=\vartheta^{\top}\bm{w}+r\|\bm{w}\|_{2}.$
Then $u_{-}\leq\varphi^{\top}\bm{w}\leq u_{+}$ and thus
$\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|=\begin{dcases}u_{-}-x,&x\leq u_{-}\\ 0,&u_{-}<x\leq u_{+}\\ x-u_{+},&x>u_{+}.\end{dcases}$
Now, we just compute separately $\int_{-\infty}^{u_{-}}$, $\int_{u_{-}}^{u_{+}}$ and $\int_{u_{+}}^{\infty}$.
$\,\int_{-\infty}^{u_{-}}\exp\left(-(1-\tau^{*})\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|\right)\mathtt{d}x$
$=$ $\,\int_{-\infty}^{u_{-}}\exp\left(-(1-\tau^{*})(u_{-}-x)\right)\mathtt{d}x$
$=$ $\,\frac{1}{1-\tau^{*}},$
$\,\int_{u_{-}}^{u_{+}}\exp\left(-(1-\tau^{*})\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|\right)\mathtt{d}x$
$=$ $\,u_{+}-u_{-}=2r\|\bm{w}\|_{2},$
and
$\,\int_{u_{+}}^{\infty}\exp\left(-(1-\tau^{*})\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|\right)\mathtt{d}x$
$=$ $\,\int_{u_{+}}^{\infty}\exp\left(-(1-\tau^{*})(x-u_{+})\right)\mathtt{d}x$
$=$ $\,\frac{1}{1-\tau^{*}}.$
Thus
$\|M_{\vartheta,r}\|_{2}^{2}=$ $\,\frac{d^{2}}{4}\|\bm{w}\|_{2}^{2}\left[\frac{2}{1-\tau^{*}}+2r\|\bm{w}\|_{2}\right]$
$=$ $\,\frac{d^{2}}{2(1-\tau^{*})}\|\bm{w}\|_{2}^{2}+\frac{d^{2}r}{2}\|\bm{w}\|_{2}^{3}.$
Since $\|\bm{w}\|_{\infty}\leq 1$, the above expression does not depend on the radius of the space $\Omega_{n}$ and hence a constant order value. ∎
Lemma 29.
Let
$f^{\tau}_{\theta}(x|\bm{w}):=\tau(1-\tau)\exp\left(-Q_{\tau}(x-\theta^{\top}\bm{w})\right),$
where $\theta\in\Omega$. If $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries, then for $\theta$ close enough to $\theta^{*}$, there exists positive constant $C$ independent of $\theta$ such that
$h(\theta^{*},\theta)\leq\phi_{n}\implies\|\theta^{*}-\theta\|_{2}\leq C\phi_{n}.$
Proof.
Recall that
$h^{2}(\theta,\theta^{*})=\frac{1}{2}\int\left(\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}\right)^{2}\mathtt{d}x$
and
$\nabla_{\theta}\sqrt{f_{\theta}^{\tau}(x|\bm{w})}=\begin{dcases}-\sqrt{\tau(1-\tau)}\frac{1-\tau}{2}\exp\left(-\frac{\tau-1}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&x\leq\theta^{\top}\bm{w}\\ \sqrt{\tau(1-\tau)}\frac{\tau}{2}\exp\left(-\frac{\tau}{2}(x-\theta^{\top}\bm{w}\right)\bm{w},&x>\theta^{\top}\bm{w}.\end{dcases}$
Consider the integrand in the following two cases:
  1. (i)
    When $x\leq\theta^{*\top}\bm{w}\leq\theta^{\top}\bm{w}$, there exists $\varphi_{1}=\theta+a_{1}(\theta^{*}-\theta)$ for some constant $a_{1}\in(0,1)$ such that
    $\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}=\sqrt{\tau(1-\tau)}\frac{\tau-1}{2}\exp\left(-\frac{\tau-1}{2}(x-\varphi_{1}^{\top}\bm{w})\right)\bm{x}^{\top}(\theta-\theta^{*}).$
    Thus
    $\,\int_{-\infty}^{\theta^{*\top}\bm{w}}\left(\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}\right)^{2}\mathtt{d}x$
    $=$ $\,\int_{-\infty}^{\theta^{*\top}\bm{w}}\frac{\tau(1-\tau)^{3}}{4}\exp\left(-(\tau-1)(x-\varphi_{1}^{\top}\bm{w})\right)(\bm{w}^{\top}(\theta-\theta^{*}))^{2}\mathtt{d}x$
    $=$ $\,\frac{\tau(1-\tau)^{2}}{4}(\bm{w}^{\top}(\theta-\theta^{*}))^{2}\exp(-(\tau-1)(\theta^{*}-\varphi_{1})^{\top}\bm{x}).$
  2. (ii)
    When $\theta^{\top}\bm{w}\leq\theta^{*\top}\bm{w}\leq x$, there exists $\varphi_{2}=\theta+a_{2}(\theta^{*}-\theta)$ for some constant $a_{2}\in(0,1)$ such that
    $\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}=\sqrt{\tau(1-\tau)}\frac{\tau}{2}\exp\left(-\frac{\tau}{2}(x-\varphi_{2}^{\top}\bm{w})\right)x^{\top}(\theta-\theta^{*}).$
    Thus
    $\,\int_{\theta^{\top}\bm{w}}^{\infty}\left(\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}\right)^{2}\mathtt{d}x$
    $=$ $\,\int_{\theta^{\top}\bm{w}}^{\infty}\frac{\tau^{3}(1-\tau)}{4}\exp\left(-\tau(x-\varphi_{2}^{\top}\bm{w})\right)(\bm{w}^{\top}(\theta-\theta^{*}))^{2}\mathtt{d}x$
    $=$ $\,\frac{\tau^{2}(1-\tau)}{4}(\bm{w}^{\top}(\theta-\theta^{*}))^{2}\exp(-\tau(\theta^{*}-\varphi_{2})^{\top}\bm{w}).$
Since the integrand in $h(\theta^{*},\theta)$ is non-negative, the integral is lower-bounded by the integral computed only on the regions described in the above two cases. When $\theta$ is close to $\theta^{*}$ and $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
$h^{2}(\theta^{*},\theta)\geq$ $\,\frac{1}{2}\frac{\tau^{*}(1-\tau^{*})^{2}}{4}(\bm{w}^{\top}(\theta-\theta^{*}))^{2}\exp(-\tau^{*}a_{3}(\theta^{*}-\theta)^{\top}\bm{w})$
$\succeq$ $\,\|\theta-\theta^{*}\|_{2}^{2}$
where $\tau^{*}=\max\{\tau,1-\tau\}$, $a_{3}=\max\{1-a_{1},1-a_{2}\}$. The last step is by applying the covariance SVD argument as in the end of Lemma 22. ∎
Appendix D Additional Discussions
D.1 For Strictly Adjacent Dataset
Definition 12.
Two datasets $\mathcal{D}$ and $\bm{X}$ are strictly adjacent if and only if $\exists\bm{x}\notin\mathcal{D}$ such that $\bm{X}=\mathcal{D}\cup\{\bm{x}\}$.
Remark 18.
Another definition for neighbouring datasets exists in the literature where one is a subset of another, and the larger dataset contains exactly one more entry.
Using the notation above, $\bm{X}=\mathcal{D}\cup\{\bm{x}\}$, so $\mathcal{D}$ and $\bm{X}$ is a pair of neighbouring datasets. In this case, (4) becomes either
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\sqrt{\eta_{n}}\left[\sqrt{\eta_{n}}+\frac{m_{n-1}(A_{n}^{c},\mathcal{D})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n-1}}(S|\mathcal{D})+\frac{m_{n-1}(A_{n}^{c},\mathcal{D})}{m_{n}(\Omega,\bm{X})}.$
or
$\mathbb{P}_{\pi_{n-1}}(S|\mathcal{D})\leq\sqrt{\eta_{n}}\left[\sqrt{\eta_{n}}+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n-1}(\Omega,\mathcal{D})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{X})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n-1}(\Omega,\mathcal{D})}.$
Corollary 30.
Under Assumptions 1-5 for any pair of strictly adjacent datasets $\bm{X}$ and $\bm{Z}$, regardless of which one is a larger set. With slight abuse of notation, let $\pi_{*}$ denote the appropriate posterior distribution $\pi_{n}$ or $\pi_{n-1}$ depending on the size of the given dataset, we have
$\mathbb{P}\Bigl{(}\mathbb{P}_{\pi_{*}}(S|\bm{X})\leq\exp(\epsilon_{n-1})\mathbb{P}_{\pi_{*}}(S|\bm{Z})+\delta_{n-1}\Bigr{)}\geq 1-Ce^{-nc_{5}\phi_{n-1}^{2}},$
where $\epsilon_{n}$ is defined with respect to $\tfrac{1}{2}\varepsilon_{n}$ and $\delta_{n}$, while $\varepsilon_{n}$,$\delta_{n}$, $t_{n}$, $c_{5}$ are defined as in Theorem 5.
Proof.
Let $n=\max\{|\bm{X}|,|\bm{Z}|\}$. Recall that the term $m_{n}(\Omega,\cdot)$ only depend on the sequence $t_{n}$ and $m_{n}(A_{n}^{c},\cdot)$ only depend on $\phi_{n}$. Denote the lagged sequences $Lt_{n}=t_{n-1}$, $L\phi_{n}=\phi_{n-1}$, and the scaled sequences $\tilde{t}_{n}=\tfrac{n}{n-1}t_{n}$, $\tilde{\phi}_{n}=\sqrt{\tfrac{n}{n-1}}\phi_{n}$. For instance, if we want to evaluate $\tfrac{m_{n-1}(A_{n}^{c},\mathcal{D})}{m_{n}(\Omega,\bm{X})}$, we can evaluate the $m_{n}(\Omega,\bm{X})$ on $Lt_{n}$ but $m_{n-1}(A_{n}^{c},\mathcal{D})$ on $\tilde{\phi}_{n-1}$, then the indices in the obtained result are synchronized to $n-1$.
Since only one sequence needs to be lagged, without loss of generality, we can always synchronize $n-1$ and thus the rest follows. ∎
D.2 Privacy in Non-identifiable Models
In the paper, we mainly addressed the case when the model is identifiable.
Definition 13 (Identifiablility).
A set of likelihood functions $f(\bm{x};\theta):\mathcal{X}\times\Omega\rightarrow\mathbb{R}_{\geq 0}$ indexed by $\theta\in\Omega$ is identifiable if and only if any likelihood function in the set is distinguishable from the rest of the set. We also call the models that adhere to that set of likelihood functions identifiable.
The ”Identifiability Assumption” does not change the level of differential privacy measured in ($\epsilon$,$\delta$). However, to determine the convergence rate for ($\epsilon$,$\delta$) in Theorem 5 requires sufficient knowledge of the model, including how the density functions are indexed.
In other words, it is not crucial to have an identifiable parameterization, in order for the model to be differentially private. In fact, under suitable regularity conditions, the level of privacy in terms of $\epsilon$ or ($\epsilon,\delta$) is intrinsic to the model, regardless of how its parameter space and likelihood function take form. The level of privacy is preserved by applying a suitable transformation (not necessarily bijective) to the parameter space and the prior. Hence, the resulting model can be considered as a reparameterization of the original model.
To see this invariance, we first consider a simple case where the parameter space is transformed by a diffeomorphism (a differentiable bijective mapping), where the change of measure can be easily deduced.
Proposition 31.
Let $(\Omega_{\psi},\Sigma_{\psi},\mathtt{d}\psi)$ and $(\Omega_{\theta},\Sigma_{\theta},\mathtt{d}\theta)$ be two measure spaces and a diffeomorphism $t:\Omega_{\psi}\rightarrow\Omega_{\theta}$ such that $t^{-1}$ is well-defined and differentiable. Then the differential privacy level of model $f_{\psi}(\bm{x};\psi)$ parameterized by $\Omega_{\psi}$ with prior $\pi_{0}(\psi)$ is the same as the privacy level of model $f_{\theta}(\bm{x};\theta)$ parameterized by $\Omega_{\theta}$ with prior $\tilde{\pi}_{0}(\theta)$,
$f_{\theta}(\bm{x};\theta)=f_{\psi}(\bm{x};t^{-1}(\theta))\cdot|J_{t^{-1}}(\theta)|,\qquad\tilde{\pi}_{0}(\theta)=\pi_{0}(t^{-1}(\psi))\cdot|J_{t^{-1}}(\theta)|$
where $J_{t^{-1}}(\theta)$ is the Jacobian of the inverse function.
Proof.
Recall that $f_{\psi}(\bm{x};\psi)$ and $\pi_{0}(\psi)$ are the likelihood and the prior with respect to the parameter space $\Omega_{\psi}$, and,
$f_{\theta}(\bm{x};\theta)=f_{\psi}(\bm{x};t^{-1}(\theta))\cdot|J_{t^{-1}}(\theta)|,\qquad\tilde{\pi}_{0}(\theta)=\pi_{0}(t^{-1}(\psi))\cdot|J_{t^{-1}}(\theta)|$
where $J_{t^{-1}}(\theta)$ is the Jacobian of the inverse function. Note that $f_{\theta}$ and $\tilde{\pi}_{0}$ are derived from $f_{\psi}$, $\pi_{0}$ through change of measure under mapping $t$.
For any quintuples $(S,\epsilon,\delta,\bm{X},\bm{Z})$, where $S\in\Sigma_{\theta}$, $\epsilon,\delta\geq 0$, $\bm{X},\bm{Z}$ are neighbouring datasets, such that
$\,\mathbb{P}_{\pi_{\theta}}(S|\bm{X})\leq e^{\epsilon}\mathbb{P}_{\pi_{\theta}}(S|\bm{Z})+\delta$
$\implies$ $\,\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{X})\leq e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{Z})+\delta$
and vice versa, since the integral is preserved through the change of measure. Therefore, a pair of $(\epsilon,\delta)$ holds for a model parameterized in $\Omega_{\psi}$ if and only if it holds for that model parameterized in $\Omega_{\theta}$ through the diffeomorphism $t$. ∎
Similarly, it is also possible to link a non-identifiable model in $\Omega_{\psi}$ to the identifiable version of itself in $\Omega_{\theta}$ while preserving the privacy level. To begin with, we need to reduce the redundancy in the non-identifiable space $\Omega_{\psi}$. Consider the following relation $\sim$ such that
$\psi_{1}\sim\psi_{2}\iff\nexists A\subset\mathcal{X},\mu(A)>0\text{ such that }\forall\bm{x}\in A,\,f(\bm{x};\psi_{1})\neq f(\bm{x};\psi_{2}).$
where $\mu$ is the base measure for the measure space $\Omega_{\psi}$. That is $\psi_{1}$ and $\psi_{2}$ lead to two likelihood functions that are equal almost everywhere. Since equal a.e. could create an issue in the DP inequality since the worst case could land on the discontinuity, we consider $f(\bm{x};\psi)$ to be continuous in $\bm{x}$. In which case, $\psi_{1}\sim\psi_{2}$ if and only if they represent the same likelihood function on the whole of $\mathcal{X}$.
Denote $\Omega_{\theta}:={{}^{\textstyle\Omega_{\psi}}/_{\textstyle\sim}}$ and consider the quotient map
$\begin{array}[]{cccc}t:&\Omega_{\psi}&\rightarrow&\Omega_{\theta}\\ &\psi&\mapsto&[\psi]\end{array}$
and if there is a way to choose every representative $[\psi]$ such that $t$ is measurable, then we have the following proposition.
Proposition 32.
Let $(\Omega_{\psi},\Sigma_{\psi},\mathtt{d}\psi)$ be a measure space such that the model with likelihood function $f_{\psi}(x;\psi)$ and prior $\pi_{0}(\psi)$ is non-identifiable. Suppose that the quotient map $t:\Omega_{\psi}\rightarrow\Omega_{\theta}\subset\Omega_{\psi}$ that reparameterizes the model onto the measure space $(\Omega_{\theta},\Sigma_{\theta},\mathtt{d}\theta)$ is measurable, then the model parameterized in $\Omega_{\theta}$ has the same level of differential privacy as the model parameterized in $\Omega_{\psi}$.
Proof.
We will prove that if $t$ is measurable,
$(\epsilon,\delta)\text{ holds for }\Omega_{\psi}\iff(\epsilon,\delta)\text{ holds for }\Omega_{\theta}$
under the same construction of the quotient map $t$, pushforward measure $t_{*}(\mathbb{P}_{\pi_{0}})$ and its density $\tilde{\pi}_{0}$ with respect to $\mathtt{d}\theta$ as in the previous proof. Since the pushforward measure and the design of the quotient ensure any integrals are preserved, the $\implies$ direction is trivial.
[…]
$\mathbb{P}_{\pi_{\theta}}(S|\bm{X})\leq e^{\epsilon}\mathbb{P}_{\pi_{\theta}}(S|\bm{Z})+\delta,$
equivalently, for any $S\in\Sigma_{\theta}$, $\bm{X},\bm{Z}$,
$\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{X})\leq e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{Z})+\delta.$
Note the following identities
$\mathbb{P}_{\pi_{\psi}}(A|\bm{X})=$ $\,\int_{A}\pi_{\psi}(\psi|\bm{X})\mathtt{d}\psi$
$=$ $\,\int_{A}\frac{\pi_{\psi}(\psi|\bm{X})}{\pi_{\psi}(\psi|\bm{Z})}\pi_{\psi}(\psi|\bm{Z})\mathtt{d}\psi$
$=$ $\,\int_{A}\left(\frac{\pi_{\psi}(\psi|\bm{X})}{\pi_{\psi}(\psi|\bm{Z})}-e^{\epsilon}\right)\pi_{\psi}(\psi|\bm{Z})\mathtt{d}\psi+e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(A|\bm{Z}).$
For $(\epsilon,\delta)$ to hold for $\Omega_{\psi}$, we need
$\int_{A}\left(\frac{\pi_{\psi}(\psi|\bm{X})}{\pi_{\psi}(\psi|\bm{Z})}-e^{\epsilon}\right)\pi_{\psi}(\psi|\bm{Z})\mathtt{d}\psi\leq\delta.$
Suppose there exists $B\subset A$ such that
$\frac{\pi_{\psi}(\psi|\bm{X})}{\pi_{\psi}(\psi|\bm{Z})}<e^{\epsilon},\quad\forall\psi\in B$
then
$\mathbb{P}_{\pi_{\psi}}(A\setminus B|\bm{X})-e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(A\setminus B|\bm{Z})\geq\mathbb{P}_{\pi_{\psi}}(A|\bm{X})-e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(A|\bm{Z}).$
Without loss of generality, it is now sufficient to check that $(\epsilon,\delta)$ hold for all sets $A\in\Sigma_{\psi}$ such that
$\frac{\pi_{\psi}(\psi|\bm{X})}{\pi_{\psi}(\psi|\bm{Z})}\geq e^{\epsilon},\quad\forall\psi\in A,$ (38)
since any other sets require a smaller $\delta$ for the differential privacy inequality to hold.
For any set $A\in\Sigma_{\psi}$ such that (38) holds. Let $S=t(A)$, note that any $\psi\in t^{-1}(S)\supseteq S$ also satisfies (38),
$\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{X})=\int_{t^{-1}(S)}\underbrace{\left(\frac{\pi_{\psi}(\psi|\bm{X})}{\pi_{\psi}(\psi|\bm{Z})}-e^{\epsilon}\right)}_{\geq 0}\pi_{\psi}(\psi|\bm{Z})\mathtt{d}\psi+e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{Z}).$
Since $A\subset t^{-1}(S)$,
$\delta\geq\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{X})-e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{Z})\geq\mathbb{P}_{\pi_{\psi}}(A|\bm{X})-e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(A|\bm{Z}).$
[…]
Remark 19.
The forward invariance in differential privacy level through mapping $t$ essentially follows from the post-processing property (Dwork et al., 2014), though note that the mapping has to be measurable for the equations to be well-defined.
{funding}
All authors were supported by the EPSRC research grant ”Pooling INference and COmbining Distributions Exactly: A Bayesian approach (PINCODE)”, reference (EP/X028100/1, EP/X028119/1, EP/X028712/1, EP/X027872/1).
LA, HD, MP and GOR were also supported by by the UKRI grant, ”On intelligenCE And Networks (OCEAN)”, reference (EP/Y014650/1).
GOR was also supported by EPSRC grants Bayes for Health (R018561), CoSInES (R034710), and EP/V009478/1.